Hugs 98

A functional programming system based on Haskell 98

User Manual

Mark P. Jones Oregon Graduate Institute

P.O. Box 91000 Portland, OR 97006

USA

John C. Peterson Department of Computer Science Yale University New Haven, CT, 06520-8285 USA

Revised version: September 10, 1999

Contents

1 Introduction 1

2 A technical summary of Hugs 98 3

3 Hugs for beginners 5 3.1 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2 Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.3 Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4 Starting Hugs 10 4.1 Environment options . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.2 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

- Set search path . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 - Set editor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 - Print statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 - Print type after evaluation . . . . . . . . . . . . . . . . . . . . . 15 - Terminate on error . . . . . . . . . . . . . . . . . . . . . . . . . 15 - Garbage collector notification . . . . . . . . . . . . . . . . . . . 16 - Literate modules . . . . . . . . . . . . . . . . . . . . . . . . . . 17 - Display dots while loading . . . . . . . . . . . . . . . . . . . . . 18 - Display nothing while loading . . . . . . . . . . . . . . . . . . . 19 - List files loaded . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 - Detailed kind errors . . . . . . . . . . . . . . . . . . . . . . . . . 19 - Use “show” to display results . . . . . . . . . . . . . . . . . . . . 20 - Import chasing . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 - Set heap size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 - Set prompt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 - Set repeat string . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 - Set preprocessor string . . . . . . . . . . . . . . . . . . . . . . . 24 - Set constraint cutoff limit . . . . . . . . . . . . . . . . . . . . . . 25

5 Hugs commands 26

5.1 Basic commands . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 - Evaluate expression . . . . . . . . . . . . . . . . . . . . . . . . . 26 - View or change settings . . . . . . . . . . . . . . . . . . . . . . . 28 - Shell escape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 - List commands . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 - Change module . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 - Change directory . . . . . . . . . . . . . . . . . . . . . . . . . . 30 - Force a garbage collection . . . . . . . . . . . . . . . . . . . . . 31 - Exit the interpreter . . . . . . . . . . . . . . . . . . . . . . . . . 31

5.2 Loading and editing modules and projects . . . . . . . . . . . . . 31 - Load definitions from module . . . . . . . . . . . . . . . . . . . 31 - Load additional files . . . . . . . . . . . . . . . . . . . . . . . . . 32 - Repeat last load command . . . . . . . . . . . . . . . . . . . . . 32 - Load project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 - Edit file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 - Find definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

5.3 Finding information about the system . . . . . . . . . . . . . . . . 35 - List names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 - Print type of expression . . . . . . . . . . . . . . . . . . . . . . . 36 - Display information about names . . . . . . . . . . . . . . . . . 36 - Display names defined in modules . . . . . . . . . . . . . . . . . 39 - Display Hugs version . . . . . . . . . . . . . . . . . . . . . . . . 39

6 Library overview 40 6.1 Standard Libraries . . . . . . . . . . . . . . . . . . . . . . . . . . 40 6.2 The Hugs-GHC Extension Libraries . . . . . . . . . . . . . . . . . 40 6.3 Portable Libraries . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 6.4 Hugs-Specific Libraries . . . . . . . . . . . . . . . . . . . . . . . . 43

7 An overview of Hugs extensions 45

7.1 Type class extensions . . . . . . . . . . . . . . . . . . . . . . . . . 46 7.1.1 Multiple parameter classes . . . . . . . . . . . . . . . . . . 46 7.1.2 More flexible instance declarations . . . . . . . . . . . . . 51 7.1.3 Overlapping instances . . . . . . . . . . . . . . . . . . . . 52 7.1.4 More flexible contexts . . . . . . . . . . . . . . . . . . . . 53

7.2 Extensible records: Trex . . . . . . . . . . . . . . . . . . . . . . . 53 7.2.1 Basic concepts . . . . . . . . . . . . . . . . . . . . . . . . . 54 7.2.2 Extensibility . . . . . . . . . . . . . . . . . . . . . . . . . . 57

7.3 Other type system extensions . . . . . . . . . . . . . . . . . . . . 60 7.3.1 Enhanced polymorphic recursion . . . . . . . . . . . . . . 60 7.3.2 Rank 2 polymorphism . . . . . . . . . . . . . . . . . . . . 60 7.3.3 Type annotations in patterns . . . . . . . . . . . . . . . . 63 7.3.4 Existential types . . . . . . . . . . . . . . . . . . . . . . . 64 7.3.5 Restricted type synonyms . . . . . . . . . . . . . . . . . . 65

7.4 Implicit parameters . . . . . . . . . . . . . . . . . . . . . . . . . . 67

8 Other Hugs programs 69 8.1 Stand-alone program execution . . . . . . . . . . . . . . . . . . . 69 8.2 Hugs for Windows . . . . . . . . . . . . . . . . . . . . . . . . . . 70

9 Conformance with Haskell 98 71 9.1 Haskell 98 features not in Hugs . . . . . . . . . . . . . . . . . . . 71 9.2 Libraries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 9.3 Haskell 98 extensions . . . . . . . . . . . . . . . . . . . . . . . . . 73

10 Pointers to further information 74

References 77

Copyright and License

The Hugs 98 system is Copyright \begin{matrix}c \\ ◯ \end{matrix} Mark P Jones, Alastair Reid, the Yale Haskell Group, and the Oregon Graduate Institute of Science and Technology, 1994-1999, All rights reserved. It is distributed as free software under the following license.

Redistribution and use in source and binary forms, with or without modification, are permitted provided that the following conditions are met:

- Redistributions of source code must retain the above copyright notice, this list of conditions and the following disclaimer.

- Redistributions in binary form must reproduce the above copyright notice, this list of conditions and the following disclaimer in the documentation and/or other materials provided with the distribution.

- Neither name of the copyright holders nor the names of its contributors may be used to endorse or promote products derived from this software without specific prior written permission.

THIS SOFTWARE IS PROVIDED BY THE COPYRIGHT HOLDERS AND THE CONTRIBUTORS ”AS IS” AND ANY EXPRESS OR IMPLIED WAR-RANTIES, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WAR-RANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE ARE DISCLAIMED. IN NO EVENT SHALL THE COPYRIGHT HOLDERS OR THE CONTRIBUTORS BE LIABLE FOR ANY DIRECT, IN-DIRECT, INCIDENTAL, SPECIAL, EXEMPLARY, OR CONSEQUENTIAL DAMAGES (INCLUDING, BUT NOT LIMITED TO, PROCUREMENT OF SUBSTITUTE GOODS OR SERVICES; LOSS OF USE, DATA, OR PROFITS; OR BUSINESS INTERRUPTION) HOWEVER CAUSED AND ON ANY THE-ORY OF LIABILITY, WHETHER IN CONTRACT, STRICT LIABILITY, OR TORT (INCLUDING NEGLIGENCE OR OTHERWISE) ARISING IN ANY WAY OUT OF THE USE OF THIS SOFTWARE, EVEN IF ADVISED OF THE POSSIBILITY OF SUCH DAMAGE.

1. Introduction

Hugs 98 is a functional programming system based on Haskell 98, the de facto standard for non-strict functional programming languages. This manual should give you all the information that you need to start using Hugs. However, it is not a tutorial on either functional programming in general or on Haskell in particular.

The first two sections provide introductory material:

• Section 2: A brief technical summary of the main features of Hugs 98, and the ways that it differs from previous releases.

• Section 3: A short tutorial on the concepts that you need to understand to be able to use Hugs.

The remaining sections provide reference material, including:

• Section 4: A summary of the command line syntax, environment variables, and command line options used by Hugs.

• Section 5: A summary of commands that can be used within the interpreter.

• Section 6: An overview of the Hugs libraries.

• Section 7: A description of Hugs extensions.

• Section 8: Information about other ways of running Hugs programs.

• Section 9: A list of differences between Hugs 98 and standard Haskell.

• Section 10: Pointers to further information.

Whether you are a beginner or a seasoned old-timer, we hope that you will enjoy working with Hugs, and that, if you will pardon the pun, you will use it to embrace functional programming!

Acknowledgements: The development of Hugs has benefited considerably from the feedback, suggestions, and bug reports provided by its users. There are too many people to name here, but thanks are due for all of their contributions. A special thank you also to our friends and colleagues at OGI, Yale, and elsewhere, for their input to the current release.

2. A technical summary of Hugs 98

Hugs 98 provides an almost complete implementation of Haskell 98 [9], including:

• Lazy evaluation, higher order functions, and pattern matching.

• A wide range of built-in types, from characters to bignums, and lists to functions, with comprehensive facilities for defining new datatypes and type synonyms.

• An advanced polymorphic type system with type and constructor class overloading.

• All of the features of the Haskell 98 expression and pattern syntax including lambda, case, conditional and let expressions, list comprehensions, do-notation, operator sections, and wildcard, irrefutable and ‘as’ patterns.

• An implementation of the main Haskell 98 primitives for monadic I/O, with support for simple interactive programs, access to text files, handle-based I/O, and exception handling.

• An almost complete implementation of the Haskell module system. The primary omission is that mutually recursive modules are not yet supported.

Hugs 98 also supports a number of advanced and experimental extensions including multi-parameter classes, extensible records, rank-2 polymorphism, existentials, scoped type variables, and restricted type synonyms. By default, these features can only be used if Hugs is started with the -98 command line flag. (See Section 7 for details.)

Hugs is implemented as an interpreter that provides:

• A relatively small, portable system that can be used on a range of different machines, from home computers, to Unix workstations.

• A read-eval-print loop for displaying the value of each expression that is entered into the interpreter.

• Fast loading, type checking, and compilation of Haskell programs, with facilities for automatic loading of imported modules.

• Integration with an external editor, chosen by the user, to allow for rapid development, and for location of errors.

• Modest browsing facilities that can be used to find information about the operations and types that are available.

Hugs is a successor to Gofer — an experimental functional programming system that was first released in September 1991 — and users of Gofer will see much that is familiar in Hugs. However, Hugs offers much greater compatibility with the Haskell standard; indeed, the name Hugs was originally chosen as a mnemonic for the “Haskell users’ Gofer system.”

There have been many modifications and enhancements to Hugs since its first release on Valentines day, February 14, in 1995. Some of the most obvious improvements include:

• Full support for new Haskell 98 features, including the labelled field syntax, do-notation, newtype, strictness annotations in datatypes, the Eval class, ISO character set, etc.

• Support for Haskell modules, and a growing collection of library modules, that includes facilities for Win32 programming.

• User interface enhancements, particularly the import chasing and search path features, which were motivated by a greater emphasis on the role of libraries in Haskell 1.3 and later versions of the language.

• Small improvements in runtime performance, and more reliable space usage, thanks to the use of non-conservative garbage collection during program execution.

• A graphical user interface for the Hugs systems that runs on the Windows operating system.

There have also been a number of other enhancements, and fixes for bugs in previous releases.

3. Hugs for beginners

This section covers the basics that you need to understand to start using Hugs. Most of the points discussed here will be familiar to anyone with experience of previous versions of Hugs or Gofer. To begin with, we need to start the interpreter; the usual way to do this is by using the command hugs, which produces a startup banner something like the following^{1}:

________________________________________________________
||   || ||  || ||  || ||__Hugs 98: Based on the Haskell 98 standard
||___|| ||__|| ||__||  __||      Copyright (c) 1994-1999
||---||         ___||            World Wide Web: http://haskell.org/hugs
||   ||                          Report bugs to: hugs-bugs@haskell.org
||   || Version: September 1999  _________________________________________
Haskell 98 mode: Restart with command line option -98 to enable extensions
Reading file "/Hugs/lib/Prelude.hs":
Hugs session for:
/Hugs/lib/Prelude.hs
Type :? for help
Prelude>

The file /Hugs/lib/Prelude.hs mentioned here contains standard definitions that are loaded into Hugs each time that the interpreter is started; the filename will vary from one installation to the next^{2}. You may notice a pause while the interpreter is initialized and the prelude definitions are loaded into the system.

3.1 Expressions

In essence, using Hugs is just like using a calculator; the interpreter simply evaluates each expression that is entered, printing the results as it goes.

Prelude> (2+3)*8
40

1On Windows 95/NT, the installation procedure normally adds Hugs to the start menu. You can also start the interpreter by double clicking on a .hs or .lhs file.

2If Hugs does not load correctly, and complains that it cannot find the prelude, then Hugs has not been installed correctly and you should look at the installation instructions.

 Prelude> sum [1..10]
 55
 Prelude>

The Prelude> characters at the begining of the first, third and fifth lines here form the Hugs prompt. This indicates that the system is ready to accept input from the user, and that it will use definitions from the Prelude module to evaluate each expression that is entered; The Hugs prelude is a special module that contains definitions for the built-in operations of Haskell, such as +, *, and sum. In response to the first prompt, the user entered the expression (2+3)*8, which was evaluated to produce the result 40. In response to the second prompt, the user typed the expression sum [1..10]. The notation [1..10] represents the list of integers between 1 and 10 inclusive, and sum is a prelude function that calculates the sum of a list of numbers. So the result obtained by Hugs is:

 1 + 2 + 3 + 4 + 5 + 6 + 7 + 8 + 9 + 10  =  55.

In fact, we could have typed this sum directly into Hugs:

 Prelude> 1 + 2 + 3 + 4 + 5 + 6 + 7 + 8 + 9 + 10
 55
 Prelude>

Unlike many calculators, however, Hugs is not limited to working with numbers; expressions can involve many different types of value, including numbers, booleans, characters, strings, lists, functions, and user-defined datatypes. Some of these are illustrated in the following example:

 Prelude> (not True) || False
 False
 Prelude> reverse "Hugs is cool"
 "looc si sguH"
 Prelude> filter even [1..10]
 [2, 4, 6, 8, 10]
 Prelude> take 10 fibs where fibs = 0:1:zipWith (+) fibs (tail fibs)
 [0, 1, 1, 2, 3, 5, 8, 13, 21, 34]
 Prelude>

You cannot create new definitions at the command prompt—these must be placed in files and loaded, as described later. The definition of fib in the last example above is local to that expression and will not be remembered for later use. Also, the expressions entered must fit on a single line.

Hugs even allows whole programs to be used as values in calculations. For example, putStr "hello, " is a simple program that outputs the string "hello, ".

Combining this with a similar program to print the string "world", gives:

 Prelude> putStr "hello, " >> putStr "world"
 hello, world
 Prelude>

Just as there are standard operations for dealing with numbers, so there are standard operations for dealing with programs. For example, the >> operator used here constructs a new program from the programs supplied as its operands, running one after the other. Normally, Hugs just prints the value of each expression entered. But, as this example shows, if the expression evaluates to a program, then Hugs will run it instead. Hugs distinguishes programs from other expressions by looking at the type of the expression entered. For example, the expression putStr "world" has type IO (), which identifies it as a program to be executed rather than a value to be printed.

3.2 Commands

Each line that you enter in response to the Hugs prompt is treated as a command to the interpreter. For example, when you enter an expression into Hugs, it is treated as a command to evaluate that expression, and to display the result. There are two commands that are particularly worth remembering:

• :q exits the interpreter. On most systems, you can also terminate Hugs by typing the end-of-file character.

• :? prints a list of all the commands, which can be useful if you forget the name of the command that you want to use.

Like most other commands in Hugs, these commands both start with a colon, :. The full set of Hugs commands is described in Section 5.

Note that the interrupt key (control-C or control-Break on most systems) can be used to abandon the process of compiling files or evaluating expressions. When the interrupt is detected, Hugs prints {Interrupted!} and returns to the prompt so that further commands can be entered.

3.3 Programs

Functions like sum, >> and take, used in the examples above, are all defined in the Hugs prelude; you can actually do quite a lot using just the types and operations provided by the prelude. But, in general, you will also want to define new types and operations, storing them in modules that can be loaded and used by Hugs. A module is simply a collection of definitions stored in a file. For example, suppose we enter the following module:

 module Fact where
 fact  :: Integer -> Integer
 fact n = product [1..n]

into a file called Fact.hs. (By convention, Hugs modules are stored in files ending with the characters .hs. The file name should match the name of the module it contains.) The product function used here is also defined in the prelude, and can be used to calculate the product of a list of numbers, just as you might use sum to calculate the corresponding sum. So the line above defines a function fact that takes an argument n and calculates its factorial. In standard mathematical notation, fact n = n!, which is usually defined by an equation:

 n! = 1 * 2 * ... * (n-1) * n

Once you become familiar with the notation, you will see that the Hugs definition is really very similar to this informal, mathematical version: the factorial of a number n is the product of the numbers from 1 to n.

Before we can use this definition in a Hugs session, we have to load Fact.hs into the interpreter. One of the simplest ways to do this uses the :load command:

 Prelude> :load fact.hs
 Reading file "fact.hs":
 Hugs session for:
 /Hugs/lib/Prelude.hs
 Fact.hs
 Fact>

Notice the list of filenames displayed after Hugs session for:; this tells you which module files are currently being used by Hugs, the first of which is always the standard prelude. The prompt is now Fact and evaluation will take place within this new module. We can start to use the fact function that we have defined:

 Fact> fact 6
 720
 Fact> fact 6 + fact 7
 5760
 Fact> fact 7 `div` fact 6
 7
 Fact>

As another example, the standard formula for the number of different ways of choosing r objects from a collection of n objects is n!/(r!(n − r)!). A simple and direct (but otherwise not particularly good) definition for this function in Hugs is as follows:

 comb n r = fact n `div` (fact r * fact (n-r))

One way to use this function is to include its definition as part of an expression entered in directly to Hugs:

 Fact> comb 5 2 where comb n r = fact n `div` (fact r * fact (n-r))
 10
 Fact>

The definition of comb here is local to this expression. If we want to use comb several times, then it would be sensible to add its definition to the file Fact.hs. Once this has been done, and the Fact.hs file has been reloaded, then we can use the comb function like any other built-in operator:

 Fact> :reload
 Reading file "fact.hs":
 Hugs session for:
 /Hugs/lib/Prelude.hs
 Fact.hs
 Fact> comb 5 2
 10
 Fact>

4. Starting Hugs

On Unix machines, the Hugs interpreter is usually started with a command line of the form:

hugs [option | file] ...

On Windows 95/NT, Hugs may be started by selecting it from the start menu or by double clicking on a file with the .hs or .lhs extension. (This manual assumes that Hugs has already been successfully installed on your system.)

Hugs uses options to set system parameters. These options are distinguished by a leading + or - and are used to customize the behaviour of the interpreter. When Hugs starts, the interpreter performs the following tasks:

• Options in the environment are processed. The variable HUGSFLAGS holds these options. On Windows 95/NT, the registry is also queried for Hugs option settings.

• Command line options are processed.

• Internal data structures are initialized. In particular, the heap is initialized, and its size is fixed at this point; if you want to run the interpreter with a heap size other than the default, then this must be specified using options on the command line, in the environment or in the registry.

• The prelude file is loaded. The interpreter will look for the prelude file on the path specified by the -P option. If the prelude, located in the file Prelude.hs, cannot be found in one of the path directories or in the current directory, then Hugs will terminate; Hugs will not run without the prelude file.

• Program files specified on the command line are loaded. The effect of a command hugs f1 ... fn is the same as starting up Hugs with the hugs command and then typing :load f1 ... fn. In particular, the interpreter will not terminate if a problem occurs while it is trying to load one of the specified files, but it will abort the attempted load command.

The environment variables and command line options used by Hugs are described in the following sections.

4.1 Environment options

Before options on the command line are processed, initial option values are set from the environment. On Windows 95/NT, these settings are added to the registry during setup. On other systems, the initial settings are determined by the HUGSFLAGS environment variable. The syntax used in this case is the same as on the command line: options are single letters, preceeded by + or -, and sometimes followed by a value. Option settings are separated by spaces; option values containing spaces are encoded using Haskell string syntax. The environment should be set up before the interpreter is used so that the search path is correctly defined to include the prelude. The built-in defaults, however, may allow Hugs to be run without any help from the environment on some systems.

It is usually more convenient to save preferred option settings in the environment rather than specifying them on the command line; they will then be used automatically each time the interpreter is started. The method for setting these options depends on the machine and operating system that you are using, and on the way that the Hugs system was installed. The following examples show some typical settings for Unix machines and PCs:

• The method for setting HUGSFLAGS on a Unix machine depends on the choice of shell. For example, a C-shell user might add something like the following to their .cshrc file:

       set HUGSFLAGS -P/usr/Hugs/lib:/usr/Hugs/libhugs -E"vi +%d %s"

The P option is used to set the search path and the E is used to set the editor. The string quotes are necessary for the value of the E option becauses it contains spaces. The setting for the path assumes that the system has been installed in /usr/local/Hugs and will need to be modified accordingly if a different directory was chosen. The editor specified here is vi, which allows the user to specify a startup line number by preceding it with a + character. The settings are easily changed to accommodate other editors.

If you are installing Hugs for the benefit of several different users, then you should probably use a script file that sets appropriate values for the environment variables, and then invokes the interpreter:

       #!/bin/sh
       HUGSFLAGS=/usr/Hugs/lib:/usr/Hugs/libhugs -E"vi +%d %s" +s
       export HUGSFLAGS
       exec /usr/local/bin/hugs $*

One advantage of this approach is that individual users do not have to worry about setting the environment variables themselves. In addition to the E and P options, other options—such as +s in this example—can be set. It is easy for more advanced users to copy and customize a script like this to suit their own needs.

• Users of DOS or Windows 3.1 might add the following line to autoexec.bat:

 set HUGSFLAGS=-P\hugs\lib;\hugs\libhugs -E"vi +%%d %%s"

The setting for the path assumes that the system has been installed in a top-level hugs directory, and will need to be modified accordingly if a different directory was chosen. In a similar way, the setting for the editor will only work if you have installed the editor program, in this case vi, that it refers to.

• On Windows 95/NT, the setup program initializes the environment, and this can be changed subsequently (on these systems only) by using either the :set command or a registry editor. The InstallShield script that performs the installation initializes the path using the installation directory; other directories can be added using -P. Installed options are stored under the HKEY_LOCAL_MACHINE key; changes to these options using :set are placed under HKEY_CURRENT_USER so that different users do not alter each other’s options.

For completeness, we should also mention the other environment variables that are used by Hugs:

• The SHELL variable on a Unix machine, or the COMSPEC variable on a DOS machine, determines which shell is used by the :! command.

• The EDITOR variable is used to try and locate an editor if no editor option has been set. Note, however, that this variable does not normally provide the extra information that is needed to be able to start the editor at a specific line in the input file.

4.2 Options

The behaviour of the interpreter, particularly the read-eval-print loop, can be customized using options. For example, you might use:

hugs -i +g +h30K

to start the interpreter with the i option (import chasing) disabled, the g option (garbage collector messages) enabled, and with a heap of thirty thousand cells.

As this example suggests, many of the options are toggles, meaning that they can either be switched on (by preceding the option with a + character) or off (by using a - character). Options may also be grouped together. For example,

hugs +stf -le is equivalent to hugs +s +t +f -l -e.

Option settings can be specified in a number of different ways—the HUGSFLAGS environment variable, the Windows registry, the command line, and the :set command—but the same syntax is used in each case. To avoid any confusion with filenames entered on the command line, option settings must always begin with a leading + or - character. However, in some cases—the h, p, r, P, and E options—the choice is not significant. With the exception of the heap size option, h, all options can be changed while the interpreter is running using the :set command. The same command can be used (without any arguments) to display a summary of the available options and to inspect their current settings.

The complete set of Hugs options is described in the sections below. The only omission here is the -98 and +98 options that are used to set the Haskell 98 compatability mode. These are discussed in Section 7.

Set search path -P⟨path⟩

The -P⟨path⟩ option changes the Hugs search path to the specified ⟨path⟩. The search path is usually initialized in the environment and should always include the directory containing the Hugs prelude and the standard libraries. When an unknown module is imported, Hugs searches for a file with the same name as the module along this path. The current directory is always searched before the path is used. Directory names should be separated by colons or, on Windows/DOS machines, by semicolons. Empty components in the path refer to the prior value of the path. For example, setting the path to dir: (dir; on Windows/DOS) would add dir to the front of the current path. Within the path, {Hugs} refers to the directory containing the Hugs libraries so one might use a path such as

{Hugs}/lib:{Hugs}/lib/hugs.

Set editor -E⟨cmd⟩

A -E⟨cmd⟩ option can be used to change the editor string to the specified ⟨cmd⟩ while the interpreter is running. The editor string is usually initialized from the environment when the interpreter starts running.

Any occurrences of %d and %s in the editor option are replaced by the start line number and the name of the file to be edited, respectively, when the editor is invoked. If specified, the line number parameter is used to let the interpreter start the editor at the line where an error was detected, or, in the case of the :find command, where a specified variable was defined.

Other editors can be selected. For example, you can use the following value to configure Hugs to use emacs:

 -E"emacs +%d %s"

More commonly, emacsclient or gnuclient is used to avoid starting a new emacs with every edit.

On Windows/DOS, you can use -Eedit for the standard DOS editor, or -Enotepad for the Windows notepad editor. However, neither edit or notepad allow you to specify a start line number, so you may prefer to install a different editor.

Print statistics +s,-s

Normally, Hugs just shows the result of evaluating each expression:

 Prelude> map (\x -> x*x) [1..10]
 [1, 4, 9, 16, 25, 36, 49, 64, 81, 100]
 Prelude> [1..]
 [1, 2, 3, 4, {Interrupted!}
 Prelude>

With the +s option, the interpreter will also display statistics about the total number of reductions and cells; the former gives a measure of the work done, while the latter gives an indication of the amount of memory used. For example:

 Prelude> :set +s
 Prelude> map (\x -> x*x) [1..10]
 [1, 4, 9, 16, 25, 36, 49, 64, 81, 100]
 (248 reductions, 429 cells)
 Prelude> [1..]
 [1, 2, 3, 4, {Interrupted!}
 (18 reductions, 54 cells)
 Prelude>

Note that the statistics produced by +s are an extremely crude measure of the behaviour of a program, and can easily be misinterpreted. For example:

• The fact that one expression requires more reductions than another does not necessarily mean that the first is slower; some reductions require much more work than others, and it may be that the average cost of reductions in the first expression is much lower than the average for the second.

• The cell count does not give any information about residency, which is the number of cells that are being used at any given time. For example, it does not distinguish between computations that run in constant space and computations with residency proportional to the size of the input.

One reasonable use of the statistics produced by +s would be to observe general trends in the behaviour of a single algorithm with variations in its input.

Print type after evaluation +t,-t

With the +t option, the interpreter will display both the result and type of each expression entered at the Hugs prompt:

 Prelude> :set +t
 Prelude> map (\x -> x*x) [1..10]
 [1, 4, 9, 16, 25, 36, 49, 64, 81, 100] :: [Int]
 Prelude> not True
 False :: Bool
 Prelude> \x -> x
 <<function>> :: a -> a
 Prelude>

Note that the interpreter will not display the type of an expression if its evaluation is interrupted or fails with a run-time error. In addition, the interpreter will not print the type, IO (), of a program in the IO monad; the interpreter treats these as a special case, giving the programmer more control over the output that is produced.

Terminate on error +f,-f

In normal use, the evaluation of an expression is abandoned completely if a runtime error occurs, such as a failed pattern match or an attempt to divide by zero. For example:

 Prelude> [1 `div` 0]
 [
 Program error: {primDivInt 1 0}
 Prelude> [1 `div` 0, 2]
 [
 Program error: {primDivInt 1 0}
 Prelude>

This is often useful during program development because it means that errors are detected as soon as they occur. However, technically speaking, the two expressions above have different meanings; the first is a singleton list, while the second has two elements. Unfortunately, the output produced by Hugs does not allow us to distinguish between the values.

The -f option can be used to make the Hugs printing option a little more accurate; this should normally be combined with -u because the built-in printer is better than the user-defined show functions at recovering from evaluation errors. With these settings, if the interpreter encounters an irreducible subexpression, then it prints the expression between a matching pair of braces and attempts to continue with the evaluation of other parts of the original expression. For the examples above, we get:

 Prelude> :set -u -f
 Prelude> [1 `div` 0]       -- value is [bottom]
 [{primDivInt 1 0}]
 Prelude> [1 `div` 0, 2]
 [{primDivInt 1 0}, 2]      -- value is [bottom, 2]
 Prelude>

Reading an expression in braces as bottom, ⊥, the output produced here shows the correct values, according to the semantics of Haskell. Of course, it is not possible to detect all occurrences of bottom like this, such as those produced by a nonterminating computation:

 Prelude> last [1..]
 ^C{Interrupted!}      -- nothing printed until interrupted
 Prelude>

Note that the basic method of evaluation is the same with both the +f and -f options; all that changes is the way that the printing mechanism deals with certain kinds of runtime error.

Garbage collector notification +g,-g

It is sometimes useful to monitor uses of the garbage collector, and to determine how many cells are recovered with each collection. If the +g option is set, then the interpreter will print a message of the form {{Gc:num}} each time that the garbage collector is invoked. The number after the colon indicates the total number of cells that are recovered.

As a simple application, we can use garbage collector messages to observe that an attempt to sum an infinite list, although non-terminating, will at least run in constant space:

 Prelude> :set +g
 Prelude> sum [1..]
 {{Gc:95763}}{{Gc:95760}}{{Gc:95760}}{{Gc:95760}}{Interrupted!}
 Prelude>

Garbage collector messages may be printed at almost any stage in a computation (or indeed while loading, type checking or compiling a file of definitions). For this reason, it is often best to turn garbage collector messages off (using :set -g, for example) if they are not required.

Literate modules +l,-l,+e,-e

Like most programming languages, Hugs usually treats source file input as a sequence of lines in which program text is the norm, and comments play a secondary role. In Hugs, as in Haskell, comments are introduced by the character sequences -- and {- ... -}.

An alternative approach, using an idea described by Knuth as “literate programming,” gives more emphasis to comments and documentation, with additional characters needed to distinguish program text from comments. Hugs supports a form of literate programming based on an idea due to Richard Bird and originally implemented as part of the functional programming language Orwell.

In a Hugs literate module, program lines are marked by a > character in the first column; any other line is treated as a program comment. This makes it particularly easy to write a document which is both an executable Hugs module and, at the same time, without need for any preprocessing, suitable for use with document preparation software such as L^{A}TEX.

Hugs will treat any input file with a name ending in .hs as a normal module and any input file with a name ending in .lhs as a literate module. If the -l option is selected, then any other file loaded into Hugs will be treated as a normal module. Conversely, if +l is selected, then these files will be treated as literate modules.

The effect of using literate modules can be thought of as applying a preprocessor to each input file that is loaded into Hugs. This has a particularly simple definition in Hugs:

 illiterate   :: String -> String
 illiterate cs = unlines [ " " ++ xs | ('>':xs) <- lines cs ]

The system of literate modules that was used in Orwell is a little more complicated than this and requires the programmer to adopt two further conventions in an attempt to catch simple errors in literate modules:

• Every input file must contain at least one line whose first character is >. This prevents modules with no definitions (because the programmer has forgotten to use the > character to mark definitions) from being accepted.

• Lines containing definitions must be separated from comment lines by one or more blank lines (i.e., lines containing only space and tab characters). This is useful for catching programs where the leading > character has been omitted from one or more lines in the definition of a function. For example:

      > map f []     = []
        map f (x:xs) = f x : map f xs

would be treated as an error.

Hugs will report on errors of this kind whenever the -e option is enabled (the default setting).

The Haskell Report defines a second style of literate programming in which code is surrounded by \begin{code} and \end{code}. See Appendix C of the Haskell Report for more information about literate programming in Haskell.

Display dots while loading +.,-.

As Hugs loads each file into the interpreter, it prints a short sequence of messages to indicate progress through the various stages of parsing the module, dependency analysis, type checking, and compilation. With the default setting, -., the interpreter prints the name of each stage, backspacing over it to erase it from the screen when the stage is complete. If you are fortunate enough to be using a fast machine, you may not always see the individual words as they flash past. After loading a file, your screen will typically look something like this:

 Prelude> :l Array
 Reading file "/Hugs/lib/Array.hs":
 Hugs session for:
 /Hugs/lib/Prelude.hs
 /Hugs/lib/Array.hs
 Prelude>

On some systems, the use of backspace characters to erase a line may not work properly—for example, if you try to run Hugs from within emacs. In this case, you may prefer to use the +. setting which prints a separate line for each stage, with a row of dots to indicate progress:

 Prelude> :load Array
 Reading file "/Hugs/lib/Array.hs":
 Parsing....................................................
 Dependency analysis........................................
 Type checking..............................................
 Compiling..................................................
 Hugs session for:
 /Hugs/lib/Prelude.hs
 /Hugs/lib/Array.hs
 Prelude>

This setting can also be useful on very slow machines where the growing line of dots provides confirmation that the interpreter is making progress through the various stages involved in loading a file. You should note, however, that the mechanisms used to display the rows of dots can add a substantial overhead to the time that it takes to load files; in one experiment, a particular program took nearly five times longer to load when the +. option was used. In this case, users might prefer to use the -q option described below.

Display nothing while loading +q,-q

The +q (quiet) option supresses the messages used to indicate progress while Hugs is loading files. If this option is turned off using the -q, then the format of output messages is determined by the current +. or -. setting.

List files loaded +w,-w

By default, Hugs prints a complete list of all the files that have been loaded into the system after every successful load or reload command. The -w option can be used to turn this feature off. Note that the :info command, without any arguments, can also be used to list the names of currently loaded files.

Detailed kind errors +k,-k

Haskell uses a system of kinds to ensure that type expressions are well-formed: for example, to make sure that each type constructor is applied to the appropriate number of arguments. For example, the following program:

 module Main where
 data Tree a  = Leaf a | Tree a :^: Tree a
 type Example = Tree Int Bool

will cause an error:

 ERROR "Main.hs" (line 3): Illegal type "Tree Int Bool" in
                           constructor application

The problem here is that Tree is a unary constructor of kind * -> *, but the definition of Example uses it as a binary constructor with at least two arguments, and hence expecting a kind of the form (* -> * -> k), for some kind k.

By default, Hugs reports problems like this with a simple message like the one shown above. However, if the +k option is selected, then the interpreter will print a more detailed version of the error message, including details about the kinds of the type expressions that are involved:

 ERROR "Main.hs" (line 3): Kind error in constructor application
 *** expression     : Tree Int Bool
 *** constructor    : Tree
 *** kind           : * -> *
 *** does not match : * -> a -> b

In addition, if the +k option is used, then Hugs will also include information about kinds in the information produced by the :info command:

 Prelude> :info Tree
 -- type constructor with kind * -> *
 data Tree a
 -- constructors:
 Leaf :: a -> Tree a
 (:^:) :: Tree a -> Tree a -> Tree a
 -- instances:
 instance Eval (Tree a)
 Prelude>

Use “show” to display results +u,-u

In normal use, Hugs displays the value of each expression entered into the interpreter by applying the standard prelude function:

 show :: Show a => a -> String

to it and displaying the resulting string of characters. This approach works well for any value whose type is an instance of the standard Show class; for example, the prelude defines instances of Show for all of the built-in datatypes. It is also easy for users to extend the class with new datatypes, either by providing a handwritten instance declaration, or by requesting an automatically derived instance as part of the datatype definition, as in:

 data Rainbow = Red | Orange | Yellow | Green | Blue | Indigo | Violet
                deriving Show

The advantage of using show is that it allows programmers to display the results of evaluations in whatever form is most convenient for users—which is not always the same as the way in which the values are represented.

This is probably all that most users will ever need. However, there are some circumstances where it is not convenient, for example, for certain kinds of debugging or for work with datatypes that are not instances of Show. In these situations, the -u option can be used to prevent the use of show. In its place, Hugs will use a built-in printing mechanism that works for all datatypes, and uses the representation of a value to determine what gets printed. At any point, the default printing mechanism can be restored by setting +u.

Import chasing +i,-i

Import chasing is a simple, but flexible mechanism for dealing with programs that involve multiple modules. It works in a natural way, using the information in import statements at the beginning of modules, and is particularly useful for large programs, or for programs that use standard Hugs libraries.

For example, consider a module Demo.hs that requires the facilities provided by the STArray library. This dependency might be reflected by including the following import statement at the beginning of Demo.hs:

 import STArray

Now, if we try to load this module into Hugs, then the system will automatically search for the STArray library and load it into Hugs, before Demo.hs is loaded. In fact, the STArray library module also begins with some import statements:

 import ST
 import Array

So, Hugs will actually load the ST and Array libraries first, then the STArray library, and only then will it try to read the rest of Demo.hs:

 Prelude> :load Demo
 Reading file "Demo.hs":

 Reading file "/hugs/libhugs/STArray.hs":
 Reading file "/hugs/libhugs/ST.hs":
 Reading file "/hugs/lib/Array.hs":
 Reading file "/hugs/libhugs/STArray.hs":
 Reading file "Demo.hs":
 Demo>

Initially, the interpreter reads only the first part of any module loaded into the system, upto and including any import statements. Only one module is allowed in each file; files with no module declaration are assumed to declare the Main module. If there are no imports, or if the modules specified as imports have already been loaded, then the system carries on and loads the module as normal. On the other hand, if the module includes import statements for modules that have not already been loaded, then the interpreter postpones the task of reading the current module until all of the specified imports have been successfully loaded. This explains why Demo.hs and STArray.hs are read twice in the example above; first to determine which imports are required, and then to read in the rest of the file once the necessary imports have been loaded.

The list of directories and filenames that Hugs tries in an attempt to locate the source for a module Mod named in an import statement can be specified by:

 [ (dir,"Mod"++suf) | dir <- [d] ++ path ++ [""],
                      suf <- ["", ".hs", ".lhs"]]

The search starts in the directory d where the file containing the import statement was found, then tries each of the directories in the current path (as defined by the -P option), represented here by path, and ends with "", which gives a search relative to the current directory. The fact that the search starts in d is particularly important because it means that you can load a multi-file program into Hugs without having to change to the directory where its source code is located. For example, suppose that /tmp contains the files, A.hs, B.hs, and C.hs, that B imports A, and that C imports B. Now, regardless of the current working directory, you can load the whole program with the command :load /tmp/C; the import in C will be taken as a reference to /tmp/B.hs, while the import in that file will be taken as a reference to /tmp/A.hs.

Import chasing is often very useful, but you should also be aware of its limitations:

• Mutually recursive modules are not supported; if A imports B, then B must not import A, either directly or indirectly through another one of its imports.

• Import chasing assumes a direct mapping from module names to the names of the files that they are stored in. If A imports B, then the code for B must be in a file called either B, B.hs, or B.lhs, and must be located in one of the directories specified above.

On rare occasions, it is useful to specify a particular pathname as the target for an import statement; Hugs allows string literals to be used as module identifiers for this purpose:

       import "../TypeChecker/Types.hs"

Note, however, that this is a nonstandard feature of Hugs, and that it is not valid Haskell syntax. You should also be aware that Hugs uses the names of files in deciding whether a particular import has already been loaded, so you should avoid situations where a single file is referred to by more than one name. For example, you should not assume that Hugs will be able to determine whether Demo.hs and ./Demo.hs are references to the same file.

Import chasing is usually enabled by default (setting +i), but it can also be disabled using the -i option.

Set heap size -h⟨size⟩

A -h⟨size⟩ option can be used to request a particular heap size for the interpreter— the total number of cells that are available at any one time—when Hugs is first loaded. The request will only be honoured if it falls within a certain range, which depends on the machine, and the version of Hugs that is used. The ⟨size⟩ parameter may include a K or k suffix, which acts as a multiplier by 1,000. For example, either of the following commands:

 hugs -h25000
 hugs -h25K

will usually start the Hugs interpreter with a heap of 25,000 cells. Cells are generally 8 bytes wide (except on the 16 bit Hugs running on DOS) and Hugs allocates a single heap. Note that the heap is used to hold an intermediate (parsed) form of each module while it is being read, type checked and compiled. It follows that, the larger the module, the larger the heap required to enable that module to be loaded into Hugs. In practice, most large programs are written (and loaded) as a number of separate modules which means that this does not usually cause problems.

Unlike all of the other options described here, the heap size setting cannot be changed from within the interpreter using a :set command. However, on Window 95/NT, changing the heap size with :set will affect the next running of Hugs since it saves all options in the registry.

Set prompt -p⟨string⟩

A -p⟨str⟩ option can be used to change the prompt to the specified string, ⟨str⟩:

 Prelude> :set -p"Hugs> "
 Hugs> :set -p"? "
 ?

Note that you will need to use quotes around the prompt string if you want to include spaces or special characters. Any %s in the prompt will be replaced by the current module name. The default prompt is "%s> ".

Set repeat string -r⟨string⟩

Hugs allows the user to recall the last expression entered into the interpreter by typing the characters $$ as part of the next expression:

 Prelude> map (1+) [1..10]
 [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]
 Prelude> filter even $$
 [2, 4, 6, 8, 10]
 Prelude>

A -r⟨str⟩ option can be used to change the repeat string—the symbol used to recall the last expression—to ⟨str⟩. For example, users of Standard ML might be more comfortable using:

 Prelude> :set -rit
 Prelude> 6 * 7
 42
 Prelude> it + it
 84
 Prelude>

Another reason to change the repeat string is to avoid clashes with uses of the same symbol in a particular program; for example, if $$ is defined as an operator in a program.

Note that the repeat string must be a valid Haskell identifier or symbol, although it will always be parsed as an identifier. If the repeat string is set to a value that is neither an identifier or symbol (for example, -r0), then the repeat last expression facility will be disabled.

Set preprocessor string -F⟨cmd⟩

Although it is not needed very often, there are sometimes applications where it useful to arrange for input files to be preprocessed before they are passed to the Hugs interpreter. This might be used, for example, to make use of filters to support conditional compilation, language extensions, literate programming systems, or format conversion. The -F option can be used to set a particular command string cmd as the name for a preprocessor. If set, then for any source file, say file.hs, that the user tries to load into Hugs, the interpreter will use the output from the command cmd file.hs instead of the contents of the file

file.hs itself.

Note that the -F option is system dependent, and is not supported on all platforms.

Set constraint cutoff limit -c⟨num⟩

The -c parameter controls the complexity of constraint satisfaction searches in the Hugs type checker. This is a technical mechanism to ensure that type checking terminates, and can safely be ignored by most users. However, in programs that make significant use of complex type class hierarchies, it may sometimes be necessary to increase the setting for the -c option to enable the Hugs type checker to explore a larger search space.

The usual default for this setting is 40, which corresponds to the command line option -c40, and we have not yet seen any examples of valid Hugs programs that are rejected with this setting. (It is possible to construct artificial programs that do require higher values, but such examples are pathological and they do not seem to appear in practice.) There is no practical benefit in choosing a lower value for this parameter. Note that, if the value used is too low, then Hugs will not be able to load some standard files and libraries, including the prelude.

5. Hugs commands

Hugs provides a number of commands that can be used to evaluate expressions, to load files, and to inspect or modify the behaviour of the system while the interpreter is running. Almost all of the commands in Hugs begin with the : character, followed by a short command word. For convenience, all but the first letter of a command may be omitted. For example, :l, :s and :q can be used as abbreviations for the :load, :set and :quit commands, respectively.

Most Hugs commands take arguments, separated from the command itself, and from one another, by spaces. The Haskell syntax for string constants can be used to enter parts of arguments that contain spaces, newlines, or other special characters. For example, the command:

 :load My File

will be treated as a command to load two files, My and File. Any of the following commands can be used to load a single file, My File, whose name includes an embedded space:

 :load "My File"
 :load "My\SPFile"
 :load "My\  \ File"
 :load My" "File

You may wish to study the lexical syntax of Haskell strings to understand some of these examples. In practice, filenames do not usually include spaces or special characters and can be entered without surrounding quotes, as in:

 :load fact.hs

The full set of Hugs commands is described in the following sections.

5.1 Basic commands

Evaluate expression ⟨expr⟩

To evaluate an expression, the user simply enters it at the Hugs prompt. This is treated as a special case, without the leading colon that is required for other commands. The expression must fit on a single line; there is no way to continue an expression onto the next line of input to the interpreter. The actual behaviour of the evaluator depends on the type of ⟨expr⟩:

• If ⟨expr⟩ has type IO t, for some type t, then it will be treated as a program using the I/O facilities provided by the Haskell IO monad. Any final result produced by the computation will be discarded.

       Prelude> putStr "Hello, world"
       Hello, world
       Prelude>

• In any other case, the value produced by the expression is converted to a string by applying the show function from the standard prelude, and the interpreter uses this to print the result.

       Prelude> "Hello" ++ ", " ++ "world"
       "Hello, world"
       Prelude>

Unlike some previous versions of Hugs, there is no special treatment for values of type String; to display a string without the enclosing quotes and special escapes, you should turn it into a program using the putStr function, as shown above.

The interpreter will not evaluate an expression that contains a syntax error, a type error, or a reference to an undefined variable:

 Prelude> sum [1..)
 ERROR: Syntax error in expression (unexpected `)')
 Prelude> sum 'a'
 ERROR: Type error in application
 *** expression     : sum 'a'
 *** term           : 'a'
 *** type           : Char
 *** does not match : [a]
 Prelude> sum [1..n]
 ERROR: Undefined variable "n"
 Prelude>

Another common problem occurs if there is no show function for the expression entered—that is, if its type is not an instance of the Show class. For example, suppose that a module defines a type T without a Show instance:

 module Test where
 data T = A | B

With just these definitions, any attempt to evaluate an expression of type T will cause an error:

 Test> A
 ERROR: Cannot find "show" function for:
 *** expression : A
 *** of type    : T
 Test>

To avoid problems like this, you will need to add an instance of the Show class to your program. One of the simplest ways to do that is to request a derived instance of Show as part of the datatype definition, as in:

 module Test where
 data T = A | B  deriving Show

Once this has been loaded, Hugs will evaluate and display values of type T:

 Test> A
 A
 Test> take 5 (cycle [A,B])
 [A, B, A, B, A]
 Test>

You should also note that the behaviour of the evaluator can be changed while the interpreter is running by using the :set command to modify option settings.

View or change settings :set [⟨options⟩]

Without any arguments, the :set command displays a list of the options and their current settings. The following output shows the settings on a typical machine:

Prelude> :set
 TOGGLES: groups begin with +/- to turn options on/off resp.
 s    Print no. reductions/cells after eval
 t    Print type after evaluation
 f    Terminate evaluation on first error
 g    Print no. cells recovered after gc
 l    Literate modules as default
 e    Warn about errors in literate modules
 .    Print dots to show progress
 q    Print nothing to show progress
 w    Always show which modules are loaded
 k    Show kind errors in full
 u    Use "show" to display results
 i    Chase imports while loading modules

 OTHER OPTIONS: (leading + or - makes no difference)
 hnum Set heap size (cannot be changed within Hugs)
 pstr Set prompt string to str
 rstr Set repeat last expression string to str
 Pstr Set search path for modules to str
 Estr Use editor setting given by str
 cnum Set constraint cutoff limit
 Fstr Set preprocessor filter to str
 Current settings: +fewkui -stgl.q -h250000 -p"%s> " -r$$ -c40
 Search path     : -P{Hugs}\lib;{Hugs}\lib\hugs;{Hugs}\lib\exts
 Editor setting  : -E"vi +%d %s"
 Preprocessor    : -F
 Compatibility   : Haskell 98 (+98)
 Prelude>

Refer to Section 4.2 for more detailed descriptions of each of these option settings.

The :set command can also be used to change options by supplying the required settings as arguments. For example:

 Prelude> :set +st
 Prelude> 1 + 3
 4 :: Int
 (4 reductions, 4 cells)
 Prelude>

On Windows 95/NT, all option settings are written out to the registry when a :set command is executed, and will be used by subsequent executions of Hugs.

Shell escape :![⟨command⟩]

A :!⟨cmd⟩ command can be used to execute the system command ⟨cmd⟩ without leaving the Hugs interpreter. For example, :!ls (or :!dir on DOS machines) can be used to list the contents of the current directory. For convenience, the :! command can be abbreviated to a single ! character.

The :! command, without any arguments, starts a new shell:

• On a Unix machine, the SHELL environment variable is used to determine which shell to use; the default is /bin/sh.

• On an DOS machine, the COMSPEC environment variable is used to determine which shell to use; this is usually COMMAND.COM.

Most shells provide an exit command to terminate the shell and return to Hugs.

List commands :?

The :? command displays the following summary of all Hugs commands:

Prelude> :?
LIST OF COMMANDS:  Any command may be abbreviated to :c where
c is the first character in the full name.
:load <filenames>   load modules from specified files
:load               clear all files except prelude
:also <filenames>   read additional modules
:reload             repeat last load command
:project <filename> use project file
:edit <filename>    edit file
:edit               edit last module
:module <module>    set module for evaluating expressions
<expr>              evaluate expression
:type <expr>        print type of expression
:?                  display this list of commands
:set <options>      set command line options
:set                help on command line options
:names [pat]        list names currently in scope
:info <names>       describe named objects
:browse <modules>   browse names defined in <modules>
:find <name>        edit module containing definition of name
:!command           shell escape
:cd dir             change directory
:gc                 force garbage collection
:version            print Hugs version
:quit               exit Hugs interpreter
Prelude>

Change module :module ⟨module⟩

A :module ⟨module⟩ command changes the current module to one given by ⟨module⟩. This is the module in which evaluation takes place and in which objects named in commands are resolved. The specified module must be part of the current program. If no module is specified, then the last module to be loaded is assumed. (Note that the name of the current module is usually displayed as part of the Hugs prompt.)

Change directory :cd ⟨directory⟩

A :cd ⟨dir⟩ command changes the current working directory to the path given by ⟨dir⟩. If no path is specified, then the command is ignored.

Force a garbage collection :gc

A :gc command can be used to force a garbage collection of the interpreter heap, and to print the number of unused cells obtained as a result:

 Prelude> :gc
 Garbage collection recovered 95766 cells
 Prelude>

Exit the interpreter :quit

The :quit command terminates the current Hugs session.

5.2 Loading and editing modules and projects

Load definitions from module :load [⟨filename⟩ ...]

The :load command removes any previously loaded modules, and then attempts to load the definitions from each of the listed files, one after the other. If one of these files contains an error, then the load process is suspended and a suitable error message will be displayed. Once the problem has been corrected, the load process can be restarted using a :reload command. On some systems, the load process will be restarted automatically after a :edit command. (The exception occurs on Windows 95/NT because of the way that the interpreter and editor are executed as independent processes.)

If no file names are specified, the :load command just removes any previously loaded definitions, leaving just the definitions provided by the prelude.

The :load command uses the list of directories specified by the current path to search for module files. We can specify the list of directory and filename pairs, in the order that they are searched, using a Haskell list comprehension:

 [ (dir,file++suf) | dir <- [""] ++ path, suf <- ["", ".hs", ".lhs"]]

The file mentioned here is the name of the module file that was entered by the user, while path is the current Hugs search path. The search starts with the directory "", which usually represents a search relative to the current working directory. So, the very first filename that the system tries to load is exactly the same filename entered by the user. However, if the named file cannot be accessed, then the system will try adding a .hs suffix, and then a .lhs suffix, and then it will repeat the process for each directory in the path, until either a suitable file has been located, or, otherwise, until all of the possible choices have been tried.

For example, this means that you do not have to type the .hs suffix to load a file Demo.hs from the current directory, provided that you do not already have a Demo file in the same directory. In the same way, it is not usually necesary to include the full pathname for one of the standard Hugs libraries. For example, provided that you do not have an Array, Array.hs, or Array.lhs file in the current working directory, you can load the standard Array library by typing

just :load Array.

Load additional files :also [⟨filename⟩ ...]

The :also command can be used to load module files, without removing any that have previously been loaded. (However, if any of the previously modules have been modified since they were last read, then they will be reloaded automatically before the additional files are read.)

If successful, a command of the form :load f1 .. fn is equivalent to the sequence of commands:

 :load
 :also f1
   .
   .
 :also fn

In particular, :also uses the same mechanisms as :load to search for modules.

Repeat last load command :reload

The :reload command can be used to repeat the last load command. If none of the previously loaded files has been modified since the last time that it was loaded, then :reload will not have any effect. However, if one of the modules has been modified, then it will be reloaded. Note that modules are loaded in a specific order, with the possibility that later modules may import earlier ones. To allow for this, if one module has been reloaded, then all subsequent modules will also be reloaded.

This feature is particularly useful in a windowing environment. If the interpreter is running in one window, then :reload can be used to force the interpreter to take account of changes made by editing modules in other windows.

Load project :project [⟨project file⟩]

Project files were originally introduced to ease the task of working with programs whose source code was spread over several files, all of which had to be loaded at the same time. The facilities for import chasing usually provide a much better way to deal with multiple file projects, but the current release of Hugs does still support the use of project files.

The :project command takes a single argument; the name of a text file containing a list of file names, separated from one another by whitespace (which may include spaces, newlines, or Haskell-style comments). For example, the following is a valid project file:

 {- A simple project file, Demo.prj -}
 Types   -- datatype definitions
 Basics  -- basic operations
 Main    -- the main program

If we load this into Hugs with a command :project Demo.prj, then the interpreter will read the project file and then try to load each of the named files. In this particular case, the overall effect is, essentially, the same as that of:

 :load Types Basics Main

Once a project file has been selected, the :project command (without any arguments) can be used to force Hugs to reread both the project file and the module files that it lists. This might be useful if, for example, the project file itself has been modified since it was first read.

Project file names may also be specified on the command line when the interpreter is invoked by preceding the project file name with a single + character. Note that there must be at least one space on each side of the +. Standard command line options can also be used at the same time, but additional filename arguments will be ignored. Starting Hugs with a command of the form hugs + Demo.prj is equivalent to starting Hugs without any arguments and then giving the command

:p Demo.prj.

The :project command uses the same mechanisms as :load to locate the files mentioned in a project file, but it will not use the current path to locate the project file itself; you must specify a full pathname.

As has already been said, import chasing usually provides a much better way to deal with multiple file programs than the old project file system. The big advantage of import chasing is that dependencies between modules are documented within individual modules, leaving the system free to determine the order in which the files should be loaded. For example, if the Main module in the example above actually needs the definitions in Types and Basics, then this will be documented by import statements, and the whole program could be loaded with a single :load Main command.

Edit file :edit [⟨file⟩]

The :edit command starts an editor program to modify or view a module file. On Windows 95/NT, the editor and interpreter are executed as independent processes. On other systems, the current Hugs session will be suspended while the editor is running. Then, when the editor terminates, the Hugs session will be resumed and any files that have been changed will be reloaded automatically. The -E option should be used to configure Hugs to your preferred choice of editor.

If no filename is specified, then Hugs uses the name of the last file that it tried to load. This allows the :edit command to integrate smoothly with the facilities for loading files.

For example, suppose that you want to load four files, f1.hs, f2.hs, f3.hs and f4.hs into the interpreter, but the file f3.hs contains an error of some kind. If you give the command:

 :load f1 f2 f3 f4

then Hugs will successfully load f1.hs and f2.hs, but will abort the load command when it encounters the error in f3.hs, printing an error message to describe the problem that occured. Now, if you use the command:

 :edit

then Hugs will start up the editor with the cursor positioned at the relevant line of f3.hs (whenever this is possible) so that the error can be corrected and the changes saved in f3.hs. When you close down the editor and return to Hugs, the interpreter will automatically attempt to reload f3.hs and then, if successful, go on to load the next file, f4.hs. So, after just two commands in Hugs, the error in f3.hs has been corrected and all four of the files listed on the original command line have been loaded into the interpreter, ready for use.

Find definition :find ⟨name⟩

The :find ⟨name⟩ command starts up the editor at the definition of a type constructor or function, specified by the argument ⟨name⟩, in one of the files currently loaded into Hugs. Note that Hugs must be configured with an appropriate editor for this to work properly. There are four possibilities:

• If there is a type constructor with the specified name, then the cursor will be positioned at the first line in the definition of that type constructor.

• If the name is defined by a function or variable binding, then the cursor will be positioned at the first line in the definition of the function or variable (ignoring any type declaration, if present).

• If the name is a constructor function or a selector function associated with a particular datatype, then the cursor will be positioned at the first line in the definition of the corresponding datatype definition.

• If the name represents an internal Hugs function, then the cursor will be positioned at the beginning of the standard prelude file.

Note that names of infix operators should be given without any enclosing them in parentheses. Thus :f !! starts an editor on the standard prelude at the first line in the definition of (!!). If a given name could be interpreted both as a type constructor and as a value constructor, then the former is assumed.

5.3 Finding information about the system

List names :names [⟨pattern⟩ ...]

The :names command can be used to list the names of variables and functions whose definitions are currently loaded into the interpreter. Without any arguments, :names produces a list of all names known to the system; the names are listed in alphabetical order.

The :names command can also accept one or more pattern strings, limiting the list of names that will be printed to those matching one or more of the given pattern strings:

 Prelude> :n fold*
 foldl foldl' foldl1 foldr foldr1
 (5 names listed)
 Prelude>

Each pattern string consists of a string of characters and may use standard wildcard syntax: * (matches anything), ? (matches any single character), \c (matches exactly the character c) and ranges of characters of the form [a-zA-Z], etc. For example:

 Prelude> :n *map* *[Ff]ile ?
 $ % * + - . / : < > appendFile map mapM mapM_readFile writeFile ^
 (17 names listed)
 Prelude>

Print type of expression :type ⟨expr⟩

The :type command can be used to print the type of an expression without evaluating it. For example:

 Prelude> :t "hello, world"
 "hello, world" :: String
 Prelude> :t putStr "hello, world"
 putStr "hello, world" :: IO ()
 Prelude> :t sum [1..10]
 sum (enumFromTo 1 10) :: (Num a, Enum a) => a
 Prelude>

Note that Hugs displays the most general type that can be inferred for each expression. For example, compare the type inferred for sum [1..10] above with the type printed by the evaluator (using :set +t):

 Prelude> :set +t
 Prelude> sum [1..10]
 55 :: Int
 Prelude>

The difference is explained by the fact that the evaluator uses the Haskell default mechanism to instantiate the type variable a in the most general type to the type Int, avoiding an error with unresolved overloading.

Display information about names :info [⟨name⟩ ...]

The :info command is useful for obtaining information about the files, classes, types and values that are currently loaded.

If there are no arguments, then :info prints a list of all the files that are currently loaded into the interpreter.

 Prelude> :info
 Hugs session for:
 /Hugs/lib/Prelude.hs
 Demo.hs
 Prelude>

If there are arguments, then Hugs treats each one as a name, and displays information about any corresponding type constructor, class, or function. The following examples show the the kind of output that you can expect:

• Datatypes: The system displays the name of the datatype, the names and types of any constructors or selectors, and a summary of related instance declarations:

Prelude> :info Either
-- type constructor
data Either a b
-- constructors:
Left :: a -> Either a b
Right :: b -> Either a b
-- instances:
instance (Eq b, Eq a) => Eq (Either a b)
instance (Ord b, Ord a) => Ord (Either a b)
instance (Read b, Read a) => Read (Either a b)
instance (Show b, Show a) => Show (Either a b)
instance Eval (Either a b)
Prelude>

Newtypes are dealt with in exactly the same way. For a simple example of a datatype with selectors, the output produced for a Time datatype:

data Time = MkTime { hours, mins, secs :: Int }

is as follows:

Time> :info Time
-- type constructor
data Time
-- constructors:
MkTime :: Int -> Int -> Int -> Time
-- selectors:
hours :: Time -> Int
mins :: Time -> Int
secs :: Time -> Int
-- instances:
instance Eval Time
Time>

• Type synonyms: The system displays the name and expansion:

Prelude> :info String
-- type constructor
type String = [Char]
Prelude>

The expansion is not included in the output if the synonym is restricted.

• Type classes: The system lists the name, superclasses, members, and instance declarations for the specified class:

Prelude> :info Num
-- type class
class (Eq a, Show a, Eval a) => Num a where
  (+) :: a -> a -> a
  (-) :: a -> a -> a
  (*) :: a -> a -> a
  negate :: a -> a
  abs :: a -> a
  signum :: a -> a
  fromInteger :: Integer -> a
  fromInt :: Int -> a
-- instances:
instance Num Int
instance Num Integer
instance Num Float
instance Num Double
instance Integral a => Num (Ratio a)
Prelude>

• Other values: For example, named functions and individual constructor, selector, and member functions are displayed with their name and type:

Time> :info . : hours min
(.) :: (a -> b) -> (c -> a) -> c -> b
(:) :: a -> [a] -> [a]  -- data constructor
hours :: Time -> Int  -- selector function
min :: Ord a => a -> a -> a  -- class member
Time>

As the last example shows, the :info command can take several arguments and prints out information about each in turn. A warning message is displayed if there are no known references to an argument:

Prelude> :info (:)
Unknown reference `(:)'
Prelude>

This illustrates that the arguments are treated as textual names for operators, not syntactic expressions (for example, identifiers). The type of the (:) operator can be obtained using the command :info : as above. There is no provision for including wildcard characters of any form in the arguments of :info commands.

If a particular argument can be interpreted as, for example, both a constructor function, and a type constructor, depending on context, then the output for both possibilities will be displayed.

Display names defined in modules :browse [⟨module⟩ ...]

The :browse command can be used to display the list of functions that are exported from the named modules:

 List> :browse Maybe
 module Maybe where
 mapMaybe :: (a -> Maybe b) -> [a] -> [b]
 catMaybes :: [Maybe a] -> [a]
 listToMaybe :: [a] -> Maybe a
 maybeToList :: Maybe a -> [a]
 fromMaybe :: a -> Maybe a -> a
 fromJust :: Maybe a -> a
 isNothing :: Maybe a -> Bool
 isJust :: Maybe a -> Bool
 List>

Only the names of currently loaded modules will be recognized.

Display Hugs version :version

The :version command is used to display the version of the Hugs interpreter:

 Prelude> :version
 -- Hugs Version September 1999
 Prelude>

This is the same information that is displayed in the Hugs startup banner.

6. Library overview

Haskell 98 places much greater emphasis on the use of libraries than early versions of the language. Following that lead, the Hugs 98 distribution includes most of the official libraries defined in the Haskell Library Report [10]. The distribution also includes a number of unofficial libraries, which fall into two categories: portable libraries, which are implemented using standard Haskell or widely implemented Haskell extensions; and Hugs-specific libraries, which use features that are not available in other Haskell implementations.

All that you need to do to use libraries is to import them using an import declaration. For example:

 module MandlebrotSet where
 import Array
 import Complex
 ...

Of course, this assumes that HUGSPATH has been set to point to the directories where the libraries are stored (Section 4.1), and that import chasing is enabled. The default search path includes the directories containing both the standard and unofficial libraries.

6.1 Standard Libraries

The Hugs 98 distribution includes the following standard libraries: Array, Char,

Complex, IO, Ix, List, Locale, Maybe, Monad, Numeric, Prelude, Random, Ratio,

and System. The libraries Directory, Time, and CPUTime, are not currently supported. The library report [10] contains full descriptions of all of theses standard libraries. Differences between the library report and the libraries supplied with Hugs are described in Section 9.

6.2 The Hugs-GHC Extension Libraries

Hugs and GHC provide a common set of libraries to aid portability; detailed specifications for these libraries are described elsewhere [12]. The Hugs-GHC modules included in the current distribution include Addr, Bits, Channel, ChannelVar,

Concurrent, Dynamic, Foreign, IOExts, Int, GetOpt, NumExts, Pretty, ST,

LazyST, Weak, and Word. The Exception and Stable libraries are not currently supported. Note that the ST and LazyST libraries cannot be used when the interpreter is running in Haskell 98 mode; the type for runST requires support for rank-2 polymorphism, which is only available in Hugs mode. (See Section 7.3.2 for further details.)

The specifications and implementations of all of these libraries are still evolving, and are subject to change.

6.3 Portable Libraries

These libraries are not part of the Haskell standard but can be ported to most Haskell systems.

• ListUtils This module provides list functions that were removed from the Prelude in the move from Haskell 1.2 to Haskell 1.3.

        module ListUtils where
        sums, products   :: Num a => [a] -> [a]
        subsequences     :: [a] -> [[a]]
        permutations     :: [a] -> [[a]]

• ParseLib This module provides a library of parser combinators, as described in the paper on Monadic Parser Combinators by Graham Hutton and Erik Meijer [5].

• Interact: This library provides facilities for writing simple interactive programs.

        module Interact where
        type Interact = String -> String
        end                :: Interact
        readChar, peekChar :: Interact -> (Char -> Interact) -> Interact
        pressAnyKey        :: Interact -> Interact
        unreadChar         :: Char -> Interact -> Interact
        writeChar          :: Char -> Interact -> Interact
        writeStr           :: String -> Interact -> Interact
        ringBell           :: Interact -> Interact
        readLine           :: String -> (String -> Interact) -> Interact

An expression e of type Interact can be executed as a program by evaluating run e.

• AnsiScreen This library defines some basic ANSI escape seqences for terminal control.

module AnsiScreen where
type Pos  = (Int,Int)
at        :: Pos -> String -> String
highlight :: String -> String
goto      :: Int -> Int -> String
home      :: String
cls       :: String

The definitions in this module will need to be adapted to work with terminals that do not support ANSI escape sequences.

• AnsiInteract This library includes both Interact and AnsiScreen, and also contains further support for screen oriented interactive I/O.

module AnsiInteract(module AnsiInteract,
                    module Interact,
                    module AnsiScreen) where
import AnsiScreen
import Interact
clearScreen       :: Interact -> Interact
writeAt           :: Pos -> String -> Interact -> Interact
moveTo            :: Pos -> Interact -> Interact
readAt            :: Pos                  -> -- start coords
                     Int                  -> -- max input length
                     (String -> Interact) -> -- continuation
                     Interact
defReadAt         :: Pos                  -> -- start coords
                     Int                  -> -- max input length
                     String               -> -- default value
                     (String -> Interact) -> -- continuation
                     Interact
promptReadAt      :: Pos                  -> -- start coords
                     Int                  -> -- max input length
                     String               -> -- prompt
                     (String -> Interact) -> -- continuation
                     Interact
defPromptReadAt   :: Pos                  -> -- start coords
                     Int                  -> -- max input length
                     String               -> -- prompt
                     String               -> -- default value
                     (String -> Interact) -> -- continuation
                     Interact

6.4 Hugs-Specific Libraries

These libraries provide several non-standard facilities for Hugs programmers. Other Haskell implementations may provide similar features, but this is not guaranteed, and there may be significant differences in organization, naming, semantics, or functionality.

• Number This library defines a numeric datatype of fixed width integers (whatever Int supplies). However, unlike the built-in Int type, overflows are detected and cause a run-time error. To ensure that all integer arithmetic in a given module includes overflow protection you must include a default declaration for Number.

  module Number where
  data Number              -- fixed width integers
  instance Eq       Number -- class instances
  instance Ord      Number
  instance Show     Number
  instance Enum     Number
  instance Num      Number
  instance Bounded  Number
  instance Real     Number
  instance Ix       Number
  instance Integral Number

This library cannot be used when Hugs is running in Haskell 98 mode because it requires features that are only supported in full Hugs mode.

• IOExtensions This module provides non-standard extensions to the IO monad.

  module IOExtensions where
  readBinaryFile     :: FilePath -> IO String
  writeBinaryFile    :: FilePath -> String -> IO ()
  appendBinaryFile   :: FilePath -> String -> IO ()
  openBinaryFile     :: FilePath -> IOMode -> IO Handle
  getCh              :: IO Char
  argv               :: [String]

• Trace: This library provides a single function, that can sometimes be useful for debugging:

  module Trace where
  trace     :: String -> a -> a
  traceShow :: Show a => String -> a -> a

When called, trace prints the string in its first argument, and then returns the second argument as its result. The traceShow function is a variant of trace that generates its output message by concatenating the supplied String argument with the result of applying show to its value argument. These functions are not referentially transparent, and should only be used for debugging, or for monitoring execution. You should also be warned that, unless you understand some of the details about the way that Hugs programs are executed, results obtained using trace can be rather confusing. For example, the messages may not appear in the order that you expect. Even ignoring the output that they produce, adding calls to trace can change the semantics of your program. Consider this a warning!

• Trex This library supports Trex extensible records. These can only be used when Hugs is compiled with Trex support using the -enable-TREX configuration option. Trex is described in more details in Section 7.2.

• HugsInternals This library provides primitives for accessing Hugs internals; for example, they provide the means with which to implement simple error-recovery and debugging facilities in Haskell. They should be regarded as an experimental feature and may not be supported in future versions of Hugs. They can only be used if hugs was configured with the

--enable-internal-prims flag.

• GenericPrint This library provides a “generic” (or “polymorphic”) print function in Haskell, that works in essentially the same way as Hugs’ builtin printer when the -u option is used. The module HugsInternals is required.

• CVHAssert This library provides a simple implementation of Cordy Hall’s assertions for performance debugging. These primitives are an experimental feature that may be removed in future versions of Hugs. They can only be used if hugs was configured with the --enable-internal-prims flag.

• Win32 This library contains Haskell versions for many of the functions in the Microsoft Win32 library. It is only available on Windows 95/NT. The --with-plugins configuration option must be used in conjunction with this and the other Microsoft libraries.

Other libraries included in the standard distribution, but not further documented

here are Sequence, Pretty, HugsDynamic, HugsLibs, StdLibs, and OldWeak.

7. An overview of Hugs extensions

The Hugs interpreter can be run in two different modes.

• Haskell 98 mode: This should be used for the highest level of compatibility with the Haskell 98 standard; known deviations from the standard are documented in Section 9. In this mode, any attempt to use Hugs specific extensions should trigger an error message. Although there are some fairly substantial differences between Haskell 1.4 and Haskell 98, our experience is that most programs written for Haskell 1.4 or earlier will need only minor modifications before they can be loaded and used from Hugs in Haskell 98 mode. Note, however, that some of the demo programs included in the standard Hugs distribution will not work in Haskell 98 mode.

• Hugs mode: This enables a number of advanced Hugs features such as type system extensions, restricted type synonyms, etc. Most of these features are described in more detail in the following sections. The underlying core language remains as in Haskell 98 mode: For example, the member function of the Functor class is still called fmap, there is no Eval class, fixity declarations can appear anywhere that a type signature is permitted, comprehension syntax is still restricted to lists, and so on.

The choice between the two modes is made when the interpreter is started, and it is (by design) not possible to change mode without exiting and restarting Hugs. The default mode is usually Haskell 98; this can also be set explicitly by starting Hugs with the command line option +98. To select the Hugs mode, you should start the interpreter with the command line option -98. The mode in which the interpreter is running is displayed as part of the startup banner, and is also included in the information produced by using the :set command without any arguments. The intention here is that beginners will get Haskell 98 mode by default, while more experienced users will be able to set up alias, batch or script files, or file associations, etc. to provide simple ways of invoking the interpreter in either mode. On Win 32 machines, for example, one can set up file associations so that you can right click on a .hs or .lhs file and get a choice of loading the file into either a Haskell 98 or Hugs mode session.

The remainder of this section sketches some of the extensions that are currently supported when the interpreter is running in Hugs mode.

7.1 Type class extensions

In Hugs mode, several of the Haskell 98 restrictions on type classes are relaxed. This allows the use of multiple parameter classes, and more flexible forms of instance declarations.

7.1.1 Multiple parameter classes

Haskell 98 allows only one type argument to be specified for any given type class. As a result, each type class corresponds to a set of types. For example, a class constraint Eq t tells us that the type t is assumed or required to be an instance of the class Eq, and the class Eq itself corresponds to the set of all equality types. In Hugs mode, this restriction is relaxed so that programmers can also define classes with multiple parameters, each of which corresponds to a multi-place relation on types.

Multiple parameter type classes seem to have many potentially interesting applications [11]. However, some practical attempts to use them have failed as a result of frustrating ambiguity problems. This occurs because the mechanisms that are used to resolve overloading are not aggressive enough. Or, to put it another way, the type relations that are defined by a collection of class and instance declarations are often too general for practical applications, where programmers might expect stronger dependencies between parameters. In the rest of this section we will describe these problems in more detail. We will also describe the mechanisms introduced in the September 1999 release of Hugs that allow programmers to declare explicit dependencies between parameters, avoiding these difficulties in many cases, and making multiple parameter classes more useful for some important practical applications.

Ambiguity problems

During the past ten years, many Haskell users have looked into the possibility of building a library for collection types, using a multiple parameter type class that looks something like the following:

   class Collects e ce where
       empty  :: ce
       insert :: e -> ce -> ce
       member :: e -> ce -> Bool

The type variable e used here represents the element type, while ce is the type of the container itself. Within this framework, we might want to define instances of this class for lists or characteristic functions (both of which can be used to represent collections of any equality type), bit sets (which can be used to represent collections of characters), or hash tables (which can be used to represent any collection whose elements have a hash function). Omitting standard implementation details, this would lead to the following declarations:

   instance Eq e => Collects e [e] where ...
   instance Eq e => Collects e (e -> Bool) where ...
   instance Collects Char BitSet where ...
   instance (Hashable e, Collects a ce)
              => Collects e (Array Int ce) where ...

All this looks quite promising; we have a class and a range of interesting implementations. Unfortunately, there are some serious problems with the class declaration. First, the empty function has an ambiguous type:

   empty :: Collects e ce => ce

By ‘ambiguous’ we mean that there is a type variable e that appears on the left of the => symbol, but not on the right. The problem with this is that, according to the theoretical foundations of Haskell overloading, we cannot guarantee a welldefined semantics for any term with an ambiguous type. For this reason, Hugs rejects any attempt to define or use such terms:

   ERROR: Ambiguous type signature in class declaration
   *** ambiguous type : Collects a b => b
   *** assigned to    : empty

We can sidestep this specific problem by removing the empty member from the class declaration. However, although the remaining members, insert and member, do not have ambiguous types, we still run into problems when we try to use them. For example, consider the following two functions:

   f x y = insert x . insert y
   g     = f True 'a'

for which Hugs infers the following types:

   f :: (Collects a c, Collects b c) => a -> b -> c -> c
   g :: (Collects Bool c, Collects Char c) => c -> c

Notice that the type for f allows the two parameters x and y to be assigned different types, even though it attempts to insert each of the two values, one after the other, into the same collection. If we’re trying to model collections that contain only one type of value, then this is clearly an inaccurate type. Worse still, the definition for g is accepted, without causing a type error. As a result, the error in this code will not be flagged at the point where it appears. Instead, it will show up only when we try to use g, which might even be in a different module.

An attempt to use constructor classes

Faced with the problems described above, some Haskell programmers might be tempted to use something like the following version of the class declaration:

   class Collects e c where
      empty  :: c e
      insert :: e -> c e -> c e
      member :: e -> c e -> Bool

The key difference here is that we abstract over the type constructor c that is used to form the collection type c e, and not over that collection type itself, represented by ce in the original class declaration. This avoids the immediate problems that we mentioned above:

• empty has type Collects e c => c e, which is not ambiguous.

• The function f from the previous section has a more accurate type:

         f :: (Collects e c) => e -> e -> c e -> c e

• The function g from the previous section is now rejected with a type error as we would hope because the type of f does not allow the two arguments to have different types.

This, then, is an example of a multiple parameter class that does actually work quite well in practice, without ambiguity problems.

There is, however, a catch. This version of the Collects class is nowhere near as general as the original class seemed to be: only one of the four instances in Section 7.1.1 can be used with this version of Collects because only one of them—the instance for lists—has a collection type that can be written in the form c e, for some type constructor c, and element type e.

Adding dependencies

To get a more useful version of the Collects class, Hugs provides a mechanism that allows programmers to specify dependencies between the parameters of a multiple parameter class^{1}.

1For readers with an interest in theoretical foundations and previous work: The use of dependency information can be seen both as a generalization of the proposal for ‘parametric

To start with an abstract example, consider a declaration such as:

   class C a b where ...

which tells us simply that C can be thought of as a binary relation on types (or type constructors, depending on the kinds of a and b). Extra clauses can be included in the definition of classes to add information about dependencies between parameters, as in the following examples:

   class D a b | a -> b where ...
   class E a b | a -> b, b -> a where ...

The notation a -> b used here between the | and where symbols—not to be confused with a function type—indicates that the a parameter uniquely determines the b parameter, and might be read as “a determines b.” Thus D is not just a relation, but actually a (partial) function. Similarly, from the two dependencies that are included in the definition of E, we can see that E represents a (partial) one-one mapping between types.

More generally, dependencies take the form x1 ... xn -> y1 ... ym, where x1, …, xn, and y1, …, yn are type variables with n> 0 and m≥ 0, meaning that the y parameters are uniquely determined by the x parameters. Spaces can be used as separators if more than one variable appears on any single side of a dependency, as in t -> a b. Note that a class may be annotated with multiple dependencies using commas as separators, as in the definition of E above. Some dependencies that we can write in this notation are redundant, and will be rejected by Hugs because they don’t serve any useful purpose, and may instead indicate an error in the program. Examples of dependencies like this include a -> a, a -> a a, a ->, etc. There can also be some redundancy if multiple dependencies are given, as in a->b, b->c, a->c, and in which some subset implies the remaining dependencies. Examples like this are not treated as errors. Note that dependencies appear only in class declarations, and not in any other part of the language. In particular, the syntax for instance declarations, class constraints, and types is completely unchanged.

By including dependencies in a class declaration, we provide a mechanism for the programmer to specify each multiple parameter class more precisely. The compiler, on the other hand, is responsible for ensuring that the set of instances that are in scope at any given point in the program is consistent with any declared dependencies. For example, the following pair of instance declarations cannot appear together in the same scope because they violate the dependency for D, even though either one on its own would be acceptable:

type classes’ that was put forward by Chen, Hudak, and Odersky [2], or as a special case of the later framework for improvement [7] of qualified types. The underlying ideas are also discussed in a more theoretical and abstract setting in a manuscript [8], where they are identified as one point in a general design space for systems of implicit parameterization.

   instance D Bool Int where ...
   instance D Bool Char where ...

Note also that the following declaration is not allowed, even by itself:

   instance D [a] b where ...

The problem here is that this instance would allow one particular choice of [a] to be associated with more than one choice for b, which contradicts the dependency specified in the definition of D. More generally, this means that, in any instance of the form:

   instance D t s where ...

for some particular types t and s, the only variables that can appear in s are the ones that appear in t, and hence, if the type t is known, then s will be uniquely determined.

The benefit of including dependency information is that it allows us to define more general multiple parameter classes, without ambiguity problems, and with the benefit of more accurate types. To illustrate this, we return to the collection class example, and annotate the original definition from Section 7.1.1 with a simple dependency:

   class Collects e ce | ce -> e where
      empty  :: ce
      insert :: e -> ce -> ce
      member :: e -> ce -> Bool

The dependency ce -> e here specifies that the type e of elements is uniquely determined by the type of the collection ce. Note that both parameters of Collects are of kind *; there are no constructor classes here. Note too that all of the instances of Collects that we gave in Section 7.1.1 can be used together with this new definition.

What about the ambiguity problems that we encountered with the original definition? The empty function still has type Collects e ce => ce, but it is no longer necessary to regard that as an ambiguous type: Although the variable e does not appear on the right of the => symbol, the dependency for class Collects tells us that it is uniquely determined by ce, which does appear on the right of the => symbol. Hence the context in which empty is used can still give enough information to determine types for both ce and e, without ambiguity. More generally, we need only regard a type as ambiguous if it contains a variable on the left of the => that is not uniquely determined (either directly or indirectly) by the variables on the right.

Dependencies also help to produce more accurate types for user defined functions, and hence to provide earlier detection of errors, and less cluttered types for programmers to work with. Recall the previous definition for a function f:

   f x y = insert x y = insert x . insert y

for which we originally obtained a type:

   f :: (Collects a c, Collects b c) => a -> b -> c -> c

Given the dependency information that we have for Collects, however, we can deduce that a and b must be equal because they both appear as the second parameter in a Collects constraint with the same first parameter c. Hence we can infer a shorter and more accurate type for f:

   f :: (Collects a c) => a -> a -> c -> c

In a similar way, the earlier definition of g will now be flagged as a type error.

Although we have given only a few examples here, it should be clear that the addition of dependency information can help to make multiple parameter classes more useful in practice, avoiding ambiguity problems, and allowing more general sets of instance declarations.

7.1.2 More flexible instance declarations

Hugs mode does not place any syntactic restrictions on the form of type expression or class constraints that can be used in an instance declaration. (Apart from the normal restrictions to ensure that such type expressions are well-formed, of course.) For example, the following definitions are all acceptable:

   instance (Eq [Tree a], Eq a) => Eq (Tree a) where ...
   instance Eq a => Eq (Bool -> a) where ...
   instance Num a => Num (String,[a]) where ...

Compare this with the restrictions of Haskell 98, which allow only variables (resp. ‘simple’ types) as the arguments of classes on the left (resp. right) hand side of the => sign. The price for this extra flexibility is that it is possible to code up arbitrarily complex instance entailments, which means that checking entailments, and hence calculating principal types, is, in the general case, undecidable. The setting for the -c option, described in Section 4.2, will cause the type checker to fail if the complexity of checking of entailments rises above a certain level. Usually, this results from examples that would otherwise cause the type checker to go into an infinite loop.

It is possible that some syntactic restrictions on instance declarations might be introduced at some point in the future in a way that will offer much of the flexibility of the current approach, but in a way that guarantees decidability.

7.1.3 Overlapping instances

The command line option +o can be used to enable support for overlapping instance declarations, provided that one of each overlapping pair is strictly more specific than the other. This facility has been introduced in a way that does not compromise the coherence of the type system. However, its semantics differs slightly from the semantics of overlapping instances in Gofer, so users may sometimes be surprised with the results. This is why we have decided to allow this feature to be turned on or off by a command line option (the default is off). If practical experience with overlapping instances is positive then we may change the current default, or even remove the option.

If the command line option +m is selected, then a lazier form of overlapping instances is supported, which we refer to as ‘multi instance resolution.’ The main idea is to omit the normal tests for overlapping instances, but to generate an error message if the type checker can find more than one way to resolve overloading for a particular instance of the class. For example, with the +m option selected, then the two instance declarations in the following program are accepted, even though they have overlapping (in fact, identical) constraints on the right of the => symbol:

   class Numeric a where describe :: a -> String
   instance Integral a => Numeric a where describe n = "Integral"
   instance Floating a => Numeric a where describe n = "Floating"

As it turns out, these instances do not cause any problems in practice because they can be distinguished by the contexts on the left of the => symbol; no standard type is an instance of both the Integral and the Floating classes:

   Main> describe (23::Int)
   "Integral"
   Main> describe (23::Float)
   "Floating"
   Main>

Note that this experimental feature may not be supported in future releases.

7.1.4 More flexible contexts

Haskell 98 allows only class constraints of the form C (a t1 ... tn) to appear in the context of any declared or inferred type, where C is a class, a is a variable, and t1, …, tn are arbitrary types (n ≥ 0). Class constraints of this form are sometimes characterized as being in head normal form. In many practical cases, we have n = 0, corresponding to class constraints of the form C a.

In Hugs mode, these restrictions are relaxed, and any type, whether in head normal form or not, is permitted to appear in a context. For example, the principal type of an expression (\x -> x==[]) is Eq [a] => [a] -> Bool, reflecting the fact that the equality function is used to compare lists of type [a]. In previous versions of Hugs, and in Haskell 98, an inferred type of Eq a => [a] -> Bool would have been produced for this term. The latter type can still be used if an explicit type signature is provided for the term, assuming that an instance declaration of the form:

    instance Eq a => Eq [a] where ...

is in scope. For example, the following program is valid:

    f  :: Eq a => [a] -> Bool
    f x = x==[]

Note that contexts are not reduced by default because this gives more general types (and potentially more efficient handling of overloading).

7.2 Extensible records: Trex

Hugs supports a flexible system of extensible records, sometimes referred to as “Trex”. The theoretical foundations for this, and a comparison with related work, is provided in a report by Gaster and Jones [3]. This section provides some background details for anybody wishing to experiment with the implementation of extensible records that is supported in the current distribution of Hugs. Please note that support for this extension in any particular build of the Hugs system is determined by a compile-time setting. If the version of Hugs that you are using was built without including support for extensible records, then you will not be able to use the features described here.

The current implementation does not use our prefered syntax for record operations; too many of the symbols that we would like to have used are already used in conflicting ways elsewhere in the syntax of Haskell 98.

7.2.1 Basic concepts

In essence, records are just collections of values, each of which is associated with a particular label. For example:

 (a = True, b = "Hello", c = 12::Int)

is a record with three components: an a field, containing a boolean value, a b field containing a string, and a c field containing the number 12. The order in which the fields are listed is not significant, so the same record value could also be written as:

 (c = 12::Int, a = True, b = "Hello")

These examples show simple ways to construct record values. We can also inspect the values held in a record using selector functions. These are written with a # character, followed immediately by the name of a field. For example:

 Prelude> #a (a = True, b = "Hello", c = 12::Int)
 True
 Prelude> #b (a = True, b = "Hello", c = 12::Int)
 "Hello"
 Prelude> #c (a = True, b = "Hello", c = 12::Int)
 12
 Prelude>

Note, howevever, that there is a conflict here with the syntax of Haskell 98 that you should be aware of if you are running in Hugs mode with an infix operator # and with support for records enabled. Under these circumstances, an expression of the form f#g will parse as f (#g) — the application of a function f to a selector function #g — and not as f # g — the application of an infix # operator to two arguments f and g. To obtain the second of these interpretations, there must be at least one space between the # and g tokens.

Record values can also be inspected by using pattern matching, with a syntax that mirrors the notation used for constructing a record. For example:

 Prelude> (\(a=x, c=y, b=_) -> (y,x)) (a = True, b = "Hello", c = 12::Int)
 (12,True)
 Prelude>

The order of fields in a record pattern is significant because it determines the order—from left to right—in which they are matched. In the following example, an attempt to match the pattern (a=[x], b=True) against the record (b=undefined, a=[]), fails because [x] does not match the empty list, but a match against (a=[2],b=True) succeeds, binding x to 2:

 Prelude> [ x | (a=[x], b=True) <- [(b=undefined, a=[]), (a=[2],b=True)]]
 [2]
 Prelude>

Changing the order of the fields in the pattern to (b=True, a=[x]) forces matching to start with the b component. But the first element in the list of records used above has undefined in its b component, so now the evaluation produces a run-time error message:

 Prelude> [ x | (b=True, a=[x]) <- [(b=undefined, a=[]), (a=[2],b=True)]]
 Program error: {undefined}
 Prelude>

Although Hugs lets you work with record values, it does not, by default, allow you to print them. More accurately, it does not automatically provide instances of the Show class for record values. So a simple attempt to print a record value will result in an error like the following:

 Prelude> (a = True, b = "Hello", c = 12::Int)
 ERROR: Cannot find "show" function for:
 *** expression : (a=True, b="Hello", c=12)
 *** of type    : Rec (a::Bool, b::[Char], c::Int)
 Prelude>

The problem here occurs because Hugs attempts to display the record by applying the show function to it, and no version of show has been defined. If you do want to be able to display record values, then you should load or import the Trex module—which is usually included in the lib/hugs directory of the Hugs distribution:

 Prelude> :load Trex
 Trex> (a = True, b = "Hello", c = 12::Int)
 (a=True, b="Hello", c=12)
 Trex> (c = 12::Int, a = True, b = "Hello")
 (a=True, b="Hello", c=12)
 Trex>

Note that the fields are always displayed with their labels in alphabetical order. The fact that the fields appear in a specific (but, frankly, arbitrary) order is very important—show is a normal function, so its output must be uniquely determined by its input, and not by the way in which that input value is written. The records used in the example above have exactly the same value, so we expect exactly the same output for each.

In a similar way, it is sometimes useful to test whether two records are equal by using the == operator. Any program that requires this feature can obtain the necessary instances of the Eq class by importing the Trex library, as shown above.

Of course, like all other values in Haskell, records have types, and these are written using expressions of the form Rec r where Rec is a built-in type constructor and r represents a ‘row’ that associates labels with types. For example:

 Trex> :t (c = 12::Int, a = True, b = "Hello")
 (a=True, b="Hello", c=12) :: Rec (a::Bool, b::[Char], c::Int)
 Trex>

The type here tells us, unsurprisingly, that the record (a=True,b="Hello",c=12) has three components: an a field containing a Bool, a b field containing a String, and a c field of type Int. As with record values themselves, the order of the components in a row is not significant:

 Trex> (a=True, b="Hello", c=12) :: Rec (b::String, c::Int, a::Bool)
 (a=True, b="Hello", c=12)
 Trex>

However, the type of a record must be an accurate reflection of the fields that appear in the corresponding value. The following example produces an error because the specified type does not list all of the fields in the record value:

 Trex> (a=True, b="Hello", c=12) :: Rec (b::String, c::Int)
 ERROR: Type error in type signature expression
 *** term           : (a=True, b="Hello", c=12)
 *** type           : Rec (a::Bool, b::[Char], c::a)
 *** does not match : Rec (b::String, c::Int)
 *** because        : field mismatch
 Trex>

Notice that Trex does not allow the kind of subtyping on record values that would allow a record like (a=True, b="Hello", c=12) to be treated implicitly as having type Rec (b::String, c::Int), simply by ‘forgetting’ about the a field. Finding an elegant, efficient, and tractable way to support this kind of implicit coercion in a way that integrates properly with other aspects of the Hugs type system remains an interesting problem for future research.

7.2.2 Extensibility

An important property of the Trex system is that the same label name can appear in many different record types, and potentially with a different value type in each case. However, all of the features that we have seen so far deal with records of some fixed ‘shape’, where the set of labels and the type of values associated with each one are fixed, and there is no apparent relationship between records of different type. In fact, all record values and record types in Trex are built-up incrementally, starting from an empty record and extending it with additional fields, one at a time. It is for this reason that Trex values are often referred to as extensible records.

In the simplest case, any given record r can be extended with a new field labelled l, provided that r does not already include an l field. For example, we can con-

struct (a=True, b="Hello") by extending (a = True) with a field b="Hello":
 Trex> (b = "Hello" | (a = True))
 (a=True, b="Hello")
 Trex>

Alternatively, we can construct the same result by extending (b = "Hello") with a field a=True:

 Trex> (a = True | (b = "Hello"))
 (a=True, b="Hello")
 Trex>

The syntax of the current implementation allows us to add several new fields at a time (the corresponding syntax for pattern matching is also supported):

 Trex> (a=True, b="Hello", c=12::Int | (b1="World"))
 (a=True, b="Hello", b1="World", c=12)
 Trex>

On the other hand, a record cannot be extended with a field of the same name, even if it has a different type. The following examples illustrate this:

 Trex> (a=True | (a=False))
 ERROR: Repeated label "a" in record (a=True, a=False)
 Trex> (a=True | r) where r = (a=12::Int)
 ERROR: (a::Int) already includes a "a" field
 Trex>

Notice that Hugs produced two different kinds of error message here. In the first case, the presence of a repeated label was detected syntactically. In the second example, the problem was detected using information about the type of the record r.

Much the same syntax can be used in patterns to decompose record values:

 Trex> (\(b=bval | r) -> (bval,r)) (a=True, b="Hello")
 ("Hello",(a=True))
 Trex>

In the previous examples, we saw how a record could be extended with new fields. As this example shows, we can use pattern matching to do the reverse operation, removing fields from a record.

We can also use pattern matching to understand how selector functions like #a, #b, and so on are implemented. For example, the selector #x is equivalent to the function (\ (x=value | _) -> value). A selector function like this is polymorphic in the sense that it can be used with any record containing an x field, regardless of the type associated with that particular component, or of any other fields that the record might contain:

 Trex> (\(x=value | _) -> value) (x=True, b="Hello")
 True
 Trex> (\(x=value | _) -> value) (name="Hugs", age=2, x="None")
 "None"
 Trex>

To understand how this works, it is useful to look at the type that Hugs assigns to this particular selector function:

 Trex> :type (\(x=value | _) -> value)
 \(x=value | _) -> value :: r\x => Rec (x::a | r) -> a
 Trex>

There are two important pieces of notation here that deserve further explanation:

• Rec (x::a | r) is the type of a record with an x component of type a. The row variable r represents the rest of the row; that is, it represents any other fields in the record apart from x. This syntax—for record type extension—was chosen to mirror the syntax that we have already seen in the examples above for record value extension.

• The constraint r\x tells us that the type on the right of the => symbol is only valid if “r lacks x,” that is, if r is a row that does not contain an x field. If you are already familiar with Haskell type classes, then you may like to think of \x as a kind of class constraint, written with postfix syntax, whose instances are precisely the rows without an x field.

For example, if we apply our selector function to a record (x=True,b="Hello") of type Rec (b::String, x::Bool), then we instantiate the variables a and r in the type above to Bool and (b::String), respectively.

In fact, the built-in selector functions have exactly the same type as the userdefined selector shown above:

 Prelude> :type #x
 #x :: b\x => Rec (x::a | b) -> a
 Prelude>

The row constraints that we see here can also occur in the type of any function that operates on record values if the types of those records are not fully determined at compile-time. For example, given the following definition:

 average r = (#x r + #y r) / 2

Hugs infers a principal type of the form:

 average :: (Fractional a, b\y, b\x) => Rec (y::a, x::a | b) -> a

However, any of the following, more specific types could be specified in a type declaration for the average function:

 average  :: (Fractional a) => Rec (x::a, y::a) -> a
 average  :: (r\x, r\y) => Rec (x::Double, y::Double | r) -> Double
 average  :: Rec (x::Double, y::Double) -> Double
 average  :: Rec (x::Double, y::Double, z::Bool) -> Double

Each of these types is an instance of the principal type given above.

These examples show an important difference between the system of records described here, and the record facilities provided by SML. In particular, SML prohibits definitions that involve records for which the complete set of fields cannot be determined at compile-time. So, the SML equivalent of the average function described above would be rejected because there is no way to determine if the record r will have any fields other than x or y. SML programmers usually avoid such problems by giving a type annotation that completely specifies the structure of the record. But, of course, if a definition is limited in this way, then it also less flexible.

With the current implementation of our type system, there is an advantage to knowing the full type of a record at compile-time because it allows the compiler to generate more efficient code. However, unlike SML, the type system also offers the extra flexibility of polymorphism and extensibility over records if that is needed.

7.3 Other type system extensions

In this section, we describe several other type system extensions that are currently available in Hugs mode.

7.3.1 Enhanced polymorphic recursion

As required by the Haskell 98 report, Hugs supports full polymorphic recursion, even for functions with overloaded types. This means that Hugs will accept definitions like the following:

 p  :: Eq a => a -> Bool
 p x = x==x && p [x]

(Note that the type signature here is not optional.) In fact, Hugs goes further than is implied by the Haskell 98 report by using programmer supplied type signatures to reduce type checking dependencies within individual binding groups. For example, the following definitions are acceptable, even though there is no explicit type signature for the function q:

 p  :: Eq a => a -> Bool
 p x = x==x && q [x]
 q x = x==x && p [x]

This is made possible by the observation that we can calculate a type for q, without needing to calculate the type of p at the same time because the type of p is already specified.

7.3.2 Rank 2 polymorphism

Hugs provides a facility that allows the definition of functions that take polymorphic arguments. This includes functions defined at the top-level, in local definitions, in class members, and in primitive declarations. In addition, Hugs allows the definition of datatypes with polymorphic and qualified types. The following examples illustrate the syntax that is used:

 amazed :: (forall a. a -> a) -> (Bool,Char)
 amazed i = (i True, i 'a')
 twice    :: (forall b. b -> f b) -> a -> f (f a)
 twice f   = f . f

There are a number of important points to note here.

• In Hugs mode, forall is a reserved word.

• Quantified variables may be of any kind, including * (types) or * -> * (unary type constructors), as in the examples above.

• Variables quantified in a forall type must appear in the scope of the quantifier. Unused quantified variables would serve no useful purpose, and are perhaps most likely to occur as the result of mispelling a variable name.

• Nested quantifiers are not allowed, and quantifiers can only appear in the types of function arguments, not in the results.

• A function can only take polymorphic arguments if an explicit type signature is provided for that function. Any call to such a function must have at least as many arguments as are needed to include the rightmost argument with a quantified type. For example, neither of the functions amazed or twice defined above can be partially applied.

• It is not necessary for all polymorphic arguments to appear at the beginning of a type signature. For example, the following type signature is valid:

 eg :: Int -> (forall a. [a] -> [a]) -> Int -> [Int]

However, as a consequence of the rules given above, the eg function defined here must always be applied to at least two arguments, even though the first of these does not have a polymorphic type.

• In the definition of a function, there must be at least as many arguments on the left hand side of the definition as are needed to included the rightmost argument with a quantified type. Only variables (or a wildcard, _) can be used as arguments on the left hand side of a function definition where a value of polymorphic type is expected.

• Arbitrary expressions can be used for polymorphic arguments in a function call, provided that they can be assigned the necessary polymorphic type. For example, all of the following expressions are valid calls to the amazed function defined above:

 amazed (let i x = x in i)
 amazed (\x -> x)
 amazed (id . id . id . id)
 amazed (id id id id id)

A similar syntax can be used to include polymorphic components in datatypes, as illustrated by the following examples:

 data Monad1 m = MkMonad1 {
                  unit1 :: (forall a. a -> m a),
                  bind1 :: (forall a b. m a -> (a -> m b) -> m b)
                 }
 data Monad2 m = MkMonad2 (forall a. a -> m a)
                          (forall a b. m a -> (a -> m b) -> m b)
 listMonad1 = MkMonad1 {unit1 = \x->[x],
                        bind1 = \x f -> concat (map f x)}
 listMonad2 = MkMonad1 (\x->[x]) (\x f -> concat (map f x))

In this case, MkMonad1 and MkMonad2 have types:

 (forall b. b -> m b) -> (forall b c. m b -> (b->m c) -> m c) -> Monad1 m
 (forall b. b -> m b) -> (forall b c. m b -> (b->m c) -> m c) -> Monad2 m

respectively, while listMonad1 and listMonad2 have types:

 Monad1 []
 Monad2 []

Note that an expression like (MkMonad2 (\x->[x])) will not be allowed because, by the rules above, the constructor MkMonad2 can only be used when both arguments are provided. An attempt to correct this problem by eta-expansion, such as (\b -> MkMonad2 (\x->[x]) b), will also fail because the new variable, b, that this introduces is now lambda-bound and hence the type that we obtain for it will not be as general as the MkMonad2 constructor requires. We can, however, use an auxiliary function with an explicit type signature to achieve the desired effect:

 halfListMonad  :: (forall a b. [a] -> (a -> [b]) -> [b]) -> Monad2 []
 halfListMonad b = MkMonad2 (\x -> [x]) b

In the current implementation, the named update syntax for Haskell datatypes (in expressions like exp{field=newValue}) cannot be used with datatypes that include polymorphic components.

The runST primitive that is used in work with lazy state threads is now handled using the facilities described here to define it as a function:

 runST :: (forall s. ST s a) -> a

As a result, it is no longer necessary to build the ST type into the interpreter; to make use of these facilities, a program should instead import the ST library (or it’s lazier variant, LazyST). A further consequence of this is that the ST and LazyST libraries cannot be used when Hugs is running in Haskell 98 mode, because that prevents the definition and use of values like runST that require rank 2 types.

7.3.3 Type annotations in patterns

Hugs allows patterns of the form (pat :: type) to be used as type annotations (in the style of Standard ML). To allow effective type inference, the type specified here must be a monotype (no forall part or class constraints are allowed), but it may include variables, which, with one exception noted below, have the same scope as the patterns in which they appear. For example, the term \(x::Int) -> x has type Int -> Int, while the expression

\(x::a) (xs::[a]) -> xs ++ [x] has type a -> [a] -> [a]. Use of this fea-

ture is subject to the following rules:

• It is an error for a variable to be used in a type where a more specific type is inferred. For example, (\(x::a) -> not x) is not a valid expression.

• It is an error for distinct variables to be used where the types concerned are the same. For example, the expression (\(x::a) (y::b) -> [x,y]) is not valid.

• Type variables bound in a pattern may be used in type signatures or further pattern type annotations within the scope of the binding. For example:

       f (x::a) = let g   :: a -> [a]
                      g y  = [x,y]
                  in  g x

In current versions of Haskell, there is no way to write a type for the local function g in this example because of the convention that free type variables are implicitly bound by a universal quantifier. In this example, the variable is instead bound in the pattern (x::a) and so the type assigned to g is actually monomorphic.

• Type signatures do not introduce bindings for type variables, but may involve type variables bound in an enclosing scope. For example, there is no direct relation between the variable t appearing in the type signature and the variable t appearing in the pattern annotation in the following code:

       pair         :: t -> s -> (t,s)
       pair x (y::t) = (x,y::t)

The explanation for this is that the type signature for pair (which might, in practice, be separated from the definition) is not in the scope of the binding of the variables x and y.

• In the current implementation, pattern type annotations that include variables are allowed on the left hand side of a pattern binding, but scope only over the right hand side of the binding.

7.3.4 Existential types

Hugs supports a form of existential types in datatype definitions in the style originally suggested by Perry and by Läufer. Existentially quantified type variables must be bound by an explicit forall construct preceding the name of the constructor in which the existentially quantified variables appear. The apparently counterintuitive use of forall to capture existentially quantified variables becomes clearer when we look at an example:

 data Appl = forall a. MkAppl (a -> Int) a (a -> a)

and consider that the MkAppl constructor defined here does indeed have a fully polymorphic type:

 MkAppl :: (a -> Int) -> a -> (a -> a) -> Appl.

Because the variable a does not appear in the result type, the choice of a in any particular use of MkAppl will be hidden. As a result, when a MkAppl constructor is used in a pattern match, we must be careful that the hidden type does not ‘escape’ into the result type or into the enclosing assumptions. For example, the following definitions are acceptable:

 good1 (MkAppl f x i) = f x
 good2 (MkAppl f x i) = map f (iterate i x)

but the next two definitions are not:

 bad1 (MkAppl f x i) = x
 bad3 y              = let g (MkAppl f x i) = length [x,y] + 1 in True

The facilities for type annotations in patterns that were described in Section 7.3.3 can be used in conjunction with existentials, as in the example:

 good (MkAppl f (x::a) i) = map f (iterate i x :: [a])

In this case, the typing annotations are redundant, although they do still provide potentially useful information for the programmer.

A datatype whose definition involves existentially quantified variables cannot use the standard Haskell mechanisms for deriving instances of standard classes like Eq and Show. If instances of these classes are required, then they must be provided explicitly by the programmer. It is possible, however, to attach type class constraints to existentially quantified variables in a datatype definition. For example, we can define a type of “show”able values using the definition:

 data Showable = forall a. Show a => MkShowable a

This will mean that all of the operations of the specified classes, in this case just Show, are available when a value of this type is unpacked during pattern matching. For example, this can be put to good use to define a simple instance of Show for the Showable datatype:

 instance Show Showable where
  show (MkShowable x) = show x

This definition can now be used in examples like the following:

 Main> map show [MkShowable 3, MkShowable True, MkShowable 'a']
 ["3", "True", "'a'"]
 Main>

7.3.5 Restricted type synonyms

Hugs supports the use of restricted type synonyms, first introduced in Gofer, and similar to the mechanisms for defining abstract datatypes that were provided in several earlier languages. The purpose of a restricted type synonym is to restrict the expansion of a type synonym to a particular set of functions. Outside of the selected group of functions, the synonym constructor behaves like a standard datatype. More precisely, a restricted type synonym definition is a top level declaration of the form:

 type T a1 ... am = rhs in f1, ..., fn

where T is a new type constructor name and rhs is a type expression typically involving some of the (distinct) type variables a1, ..., am. The major difference with a normal type synonym definition is that the expansion of the type synonym can only be used within the binding group of one of the functions f1, …, fn (all of which must be defined by top-level definitions in the module containing the restricted type synonym definition). In the definition of any other value, T is treated as if it had been introduced by a definition of the form:

 data T a1 ... am = ...

For a simple example of this, consider the following definition of a datatype of stacks in terms of the standard list type:

 type Stack a = [a] in emptyStack, push, pop, top, isEmpty
 emptyStack :: Stack a
 emptyStack  = []
 push       :: a -> Stack a -> Stack a
 push        = (:)

 pop        :: Stack a -> Stack a
 pop []      = error "pop: empty stack"
 pop (_:xs)  = xs
 top        :: Stack a -> a
 top []      = error "top: empty stack"
 top (x:_)   = x
 isEmpty    :: Stack a -> Bool
 isEmpty     = null

The type signatures here are particularly important. For example, because emptyStack is mentioned in the definition of the restricted type synonym Stack, the definition of emptyStack is type correct. The declared type for emptyStack is Stack a which can be expanded to [a], agreeing with the type for the empty list []. However, in an expression outside the binding group of these functions, the Stack a type is quite distinct from the [a] type:

 ? emptyStack ++ [1]
 ERROR: Type error in application
 *** Expression     : emptyStack ++ [1]
 *** Term           : emptyStack
 *** Type           : Stack b
 *** Does not match : [a]
 ?

The binding group of a value is to the set of values whose definitions are in the same mutually recursive group of bindings. In particular, this does not extend to class and instance declarations so we can define instances such as:

 instance Eq a => Eq (Stack a) where
     s1 == s2 | isEmpty s1 = isEmpty s2
              | isEmpty s2 = isEmpty s1
              | otherwise  = top s1 == top s2 && pop s1 == pop s2

As a convenience, Hugs allows the type signatures of functions mentioned in the type synonym declaration to be specified within the definition. Thus the above example could also have been written as:

 type Stack a = [a] in
     emptyStack :: Stack a,
     push       :: a -> Stack a -> Stack a,
     pop        :: Stack a -> Stack a,
     top        :: Stack a -> a,
     isEmpty    :: Stack a -> Bool
 emptyStack  = []
 ...

If a type signature is included as part of the definition of a restricted type synonym, then the declaration should not be repeated elsewhere in the module; Hugs will reject any attempt to do this by complaining about a repeated type signature.

7.4 Implicit parameters

Hugs supports an experimental implementation of Implicit Parameters, which provides a technique for introducing dynamic binding of variables into a language with a Hindley-Milner based type system. This is based on as-yet-unpublished work by Jeff Lewis, Erik Meijer and Mark Shields. The prototype implementation, and much of the following description, was provided by Jeff Lewis.

A variable is called dynamically bound when it is bound by the calling context of a function and statically bound when bound by the callee’s context. In Haskell, all variables are statically bound. Dynamic binding of variables is a notion that goes back to Lisp, but was later discarded in more modern incarnations, such as Scheme. Dynamic binding can be very confusing in an untyped language, and unfortunately, typed languages, in particular Hindley-Milner typed languages like Haskell, only support static scoping of variables.

However, by a simple extension to the type class system of Haskell, we can support dynamic binding. Basically, we express the use of a dynamically bound variable as a constraint on the type. These constraints lead to types of the form (?x::t') => t, which says “this function uses a dynamically-bound variable ?x of type t'”. For example, the following expresses the type of a sort function, implicitly parameterized by a comparison function named cmp.

 sort :: (?cmp :: a -> a -> Bool) => [a] -> [a]

The dynamic binding constraints are just a new form of predicate in the type class system.

An implicit parameter is introduced by the special form ?x, where x is any valid identifier. Use if this construct also introduces new dynamic binding constraints. For example, the following definition shows how we can define an implicitly parameterized sort function in terms of an explicitly parameterized sortBy function:

 sortBy :: (a -> a -> Bool) -> [a] -> [a]
 sort   :: (?cmp :: a -> a -> Bool) => [a] -> [a]
 sort    = sortBy ?cmp

Dynamic binding constraints behave just like other type class constraints in that they are automatically propagated. Thus, when a function is used, its implicit parameters are inherited by the function that called it. For example, our sort function might be used to pick out the least value in a list:

 least   :: (?cmp :: a -> a -> Bool) => [a] -> a
 least xs = fst (sort xs)

Without lifting a finger, the ?cmp parameter is propagated to become a parameter of least as well. With explicit parameters, the default is that parameters must always be explicit propagated. With implicit parameters, the default is to always propagate them.

However, an implicit parameter differs from other type class constraints in the following way: All uses of a particular implicit parameter must have the same type. This means that the type of (?x, ?x) is (?x::a) => (a, a), and not (?x::a, ?x::b) => (a, b), as would be the case for type class constraints.

An implicit parameter is bound using an expression of the form e with binds, or equivalently as dlet binds in e, where both with and dlet (dynamic let) are new keywords. These forms bind the implicit parameters arising in the body, not the free variables as a let or where would do. For example, we define the min function by binding cmp.

 min :: [a] -> a
 min  = least with ?cmp = (<=)

Syntactically, the binds part of a with or dlet construct must be a collection of simple bindings to variables (no function-style bindings, and no type signatures); these bindings are neither polymorphic or recursive.

8. Other Hugs programs

The Hugs interpreter is available in two other guises: a stand-alone system that executes programs in a ‘load and go’ style, without the surrounding command system; and a Windows user interface, layered on top of the basic Hugs system.

8.1 Stand-alone program execution

Once a program has been developed and debugged, the Hugs command loop can be eliminated and the program can be executed immediately without any command to run it. A slightly modified version of the interpreter called runhugs loads the literate program specified as its first argument and runs main in module Main. Unlike the standard Hugs system, runhugs makes command arguments available to the running Hugs system. The first argument, specifying the program, is removed from the argument list.

On Unix systems, executable programs may be created by placing runhugs in the first line of an executable file, like so:

#!/hugs/runhugs
> module Main where
> main = putStr "Hello, World\n"

Because runHugs uses literate Haskell only, the line starting with #! is viewed as a comment. Stand-alone programs can import other modules using import chasing—these modules need not be literate. The runhugs program uses the same environment variables to set Hugs options as the standard Hugs systems. However, runhugs does not set options from the command line; all command line options are passed into the executing Hugs program. The stand-alone Hugs program may return an exit code.

On Windows 95/NT, runhugs is invoked using a separate file extension that is set up to call runhugs rather than hugs. Installation sets up the .hsx extension for this purpose. A .hsx program will run when it is clicked on; a console window will appear if the program writes to standard output or reads from standard input. This window is closed immedately upon exiting the program. There is no way to pass parameters to the .hsx program when it is double-clicked. Windows 95/NT can also use runhugs to open files of a given type; this involves setting the “open” command for the file type to call runhugs, passing it the Haskell program to run and the file being opened. The online documentation has some examples of this.

8.2 Hugs for Windows

Hugs for Windows (winhugs) offers a GUI front-end to the Hugs interpreter on Microsoft Windows platforms. The user interface features a scrolling console window that mimics the normal Hugs interface, together with a menu and toolbar that provide additional facilities for browsing Haskell programs. Most of the additional features are self-explanatory, although short descriptions of menu and toolbar choices are displayed in a status line. Hugs for Windows uses the same command line options and environment/registry variables as Hugs. It also stores options in a .ini file.

The Hugs for Windows front-end is useful for beginners, but is not compatible with the Win32 libraries or with programs that use them, such as Conal Elliot’s Fran system or Paul Hudak’s Graphics library. In addition, the current implementation uses a compute-intensive polling process to detect certain events, and this can incur a fairly substantial performance penalty. For these reasons, the Hugs for Windows front-end is not recommended for work on large projects.

9. Conformance with Haskell 98

A number of Haskell 98 features are not yet implemented in Hugs 98. All known differences between the specification and implementation are described in this section, although there are bound to be some unintentional omissions.

9.1 Haskell 98 features not in Hugs

• Mutually recursive modules are not supported.

• Some library functions have been moved into the Prelude. This is necessary because the Prelude and the standard libraries, as defined in the Haskell 98 report, are mutually recursive. This mutual recursion has been avoided by moving the following functions into the Prelude:

  – From Ix: Ix(range, index, inRange, rangeSize).
  – From Char: isAscii, isControl, isPrint, isSpace, isUpper, isLower,
    isAlpha, isDigit, isOctDigit, isHexDigit, isAlphanum, digitToInt,
    intToDigit, toUpper, toLower, ord, and chr.
  – From Ratio: Ratio, Rational, (%), numerator, denominator, and
    approxRational.

• Derived Read instances do not work for some infix constructors. If an infix constructor has left associativity and the type appears recursively on the left side of the constructor, then the read instance will loop.

• Hugs does not allow the use of qualified names in instance declarations.

• Hugs does not use the Unicode character set yet. Characters are currently drawn from the ISO Latin-1 set.

• Two adjacent dashes -- start a one line comment; for strictly technical reasons, the change from Haskell 1.4 to Haskell 98 to use maximal munch for such comments has not yet been implemented.

• The floating point printer is not exactly as defined in the report. The printed form of a floating point number may re-read as a slightly different number.

• Derived instances for large tuples are not supplied. Instances for tuples larger than 5 (3 in the 16 bit PC system) are not in the Prelude.

• When using getArgs, only the stand-alone system passes arguments to the executing program. The interactive system always uses an empty argument list when runnning a program.

• The syntax of sections is slightly different. For example, the Haskell expression (2*3+) must instead be written as ((2*3)+).

• Instead of IO.hIsEOF, hugs provides IO.hugsHIsEOF. Whereas hIsEOF should tell you if the next call of hGetChar would raise an EOF error; hugsHIsEOF tells you if the last call of hGetCHar raised an EOF error (the same as ANSI C’s feof).

• We ignore entity lists in qualified imports (but unqualified imports are treated correctly). For example, you can write:

    import qualified Prelude ( foo )

even though foo is not exported from the Prelude and you can write:

  module M() where
  import qualified Prelude () -- import nothing
  x = Prelude.length "abcd"

• The Double type is implemented as a single precision float (this isn’t forbidden by the standard but it is unusual).

9.2 Libraries

The following libraries are not yet available: Directory, Time, CPUTime, Bit,

Nat, and Signed.

In the IO library, these functions are not defined: handlePosn, ReadWriteMode

hFileSize, hIsEOF, isEOF, hSetBuffering, hGetBuffering, hSeek, hIsSeekable,

hReady, and hLookahead. The following non-standard functions are exported:

hugsGetCh :: IO Char   -- getchar without echoing to screen
hugsHIsEOF :: Handle -> IO Bool
-- same semantics as C's "feof" (different from Haskell's hIsEOF)
hugsIsEOF  :: IO Bool
-- same semantics as C's "feof(stdin)"
hPutStrLn  :: String -> IO ()
-- corresponds to Prelude.putStrLn

9.3 Haskell 98 extensions

In addition to the features described in Section 7, Hugs 98 supports some modest extensions to the Haskell language.

• Import declarations may specify a file name instead of a module name.

• The T(..) syntax is allowed for type synonyms in import and export lists.

10. Pointers to further information

Hugs

The full distribution for Hugs is available on the World Wide Web from:

     http://haskell.org/hugs.

The distribution includes source code, demo programs, library files, user documentation, and precompiled binaries for common platforms.

There is a mailing list for Hugs users at hugs-users@haskell.org, and another for bug reports at hugs-bugs@haskell.org. Admin requests (for example, to subscribe or unsubscribe) should be sent to majordomo@haskell.org. For more detailed instructions, just send a message to this address with help in the body. An overview of nearly all Haskell related resources can be found at

     http://haskell.org.

Functional programming

The usenet newsgroup comp.lang.functional provides a forum for general discussion about functional programming languages. A list of frequently asked questions (FAQs), and their answers, is available from:

     http://www.cs.nott.ac.uk/Department/Staff/gmh/faq.html.

The FAQ list contains many pointers to other functional programming resources around the world.

Further reading

As we said at the very beginning, this manual is not intended as a tutorial on either functional programming in general, or Haskell in particular. For these things, our first recommendations would be for the Introduction to Functional Programming by Bird and Wadler [1], and the Gentle Introduction to Haskell by Hudak, Peterson and Fasel [4], respectively. Note, however, that there are several other good textbooks dealing either with Haskell or related languages.

For those with an interest in the implementation of Hugs, the report about the implementation of Gofer [6], Hugs’ predecessor, should be a useful starting point.

Bibliography

[1] R. Bird and P. Wadler. Introduction to functional programming. Prentice Hall, 1988.

[2] K. Chen, P. Hudak, and M. Odersky. Parametric type classes (extended abstract). In ACM conference on LISP and Functional Programming, San Francisco, CA, June 1992.

[3] B. R. Gaster and M. P. Jones. A polymorphic type system for extensible records and variants. Technical Report NOTTCS-TR-96-3, Computer Science, University of Nottingham, November 1996.

[4] P. Hudak and J. Fasel. A gentle introduction to Haskell. ACM SIGPLAN Notices, 27(5), May 1992. Also available as Research Report YALEU/DCS/RR-901, Yale University, Department of Computer Science, April 1992.

[5] G. Hutton and E. Meijer. Monadic parser combinators. Available from

http://www.cs.nott.ac.uk/Department/Staff/gmh/bib.html, 1996.

[6] M. Jones. The implementation of the Gofer functional programming system. Research Report YALEU/DCS/RR-1030, Yale University, New Haven, Connecticut, USA, May 1994. Available on the World-Wide Web from

http://www.cse.ogi.edu/~mpj/pubs.html.

[7] M. P. Jones. Simplifying and improving qualified types. In International Conference on Functional Programming Languages and Computer Architecture, pages 160–169, June 1995.

[8] M. P. Jones. Exploring the design space for type-based implicit parameterization. July 1999.

[9] S. Peyton Jones and J. Hughes (editors). Report on the Programming Language Haskell 98, A Non-strict Purely Functional Language, February 1999.

Available from http://www.haskell.org/definition/.

[10] S. Peyton Jones and J. Hughes (editors). Standard libraries for the Haskell 98 programming language, February 1999. Available from

http://www.haskell.org/definition/.

[11] S. Peyton Jones, M. Jones, and E. Meijer. Type classes: Exploring the design space. In Proceedings of the Second Haskell Workshop, Amsterdam, June 1997. Available on the web from

http://www.cse.ogi.edu/~mpj/pubs/multi.html.

[12] The Hugs/GHC Team. The Hugs-GHC Extension Libraries, January 1999.

Available from http://www.haskell.org/libraries/.

Index

options

+.,-., 18
+e,-e, 17
+f,-f, 15
+g,-g, 16
+i,-i, 21
+k,-k, 19
+l,-l, 17
+q,-q, 19
+s,-s, 14
+t,-t, 15

+u,-u, 16, 20

+w,-w, 19

-F, 24 -P, 13 -c, 25 -h, 23 -p, 24 -r, 24

commands, 7, 26 :!, 29 :?, 7, 30

:also, 32
:browse, 39
:cd, 30

:edit, 31, 34

:find, 34
:gc, 31
:info, 36

:load, 31–34

:module, 30
:names, 35
:project, 32
:quit, 7, 31
:reload, 31, 32

:set, 13, 23, 28

:type, 36

                                   :version, 39

emacs editor, 14, 18

evaluator, 26

garbage collection, 4, 16, 31 Gofer, 4, 75

Haskell, 3 Haskell98, 1 heap size, 10, 13, 23

                                HUGSPATH, 40

import chasing, 4, 21, 33, 40 interrupt key, 7 IO monad, 3, 15, 27

libraries, 40

                                   AnsiInteract, 42
                                   AnsiScreen, 42

Array, 21, 32

                                   IOExtensions, 43
                                   Interact, 41
                                   ListUtils, 41
                                   Number, 43
                                   ParseLib, 41
                                   STArray, 21

ST, 21

                                   Trace, 43
                                   Trex, 44
                                   Win32, 44

modules, 17, 31 literate, 17

options, 12, 28, 33

                                   +t,-t, 36

-E, 13 -P, 10 -h, 13

prelude, 5, 7, 8, 10, 27, 31, 35 prompt, 6, 24, 26 vi editor, 11